### v1

# Beyond Sociodemographic Prompting: Using Supervision to Align LLMs with Human Response Distributions

Abstract
The ability to accurately predict how different population groups would answer subjective questions would have great value. In this work, we show that use of relatively simple supervision can greatly improve language model alignment with diverse population groups, as measured over three datasets spanning various topics. Beyond evaluating average performance, we also report how alignment varies across specific groups. The simplicity and generality of our approach promotes easy adoption, while our broad findings provide useful guidance for when to use or not use our approach in practice. By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a useful benchmark to stimulate future research 111Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment.
1 Introduction
Human surveys elicit valuable public opinion on a wide range of topics, such as proposed legislation, candidates for office (i.e., election polls), marketing or outreach campaigns, commercial products and services, etc. Hayati et al. (2024); Santurkar et al. (2023). Subjective annotation tasks in NLP also seek diverse, human judgments Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). Given the frequency, breadth, and importance of obtaining opinion data from diverse populations, there has been great interest in developing effective ways to conduct such surveys accurately and efficiently.
Recent work has sought to simulate human responses using LLMs in many areas, such as in psychology, sociology, and economic studies Aher et al. (2023), NLP annotation Nasution and Onan (2024), and large-scale survey creation and testing Rothschild et al. (2024). More specifically, LLMs are being increasingly used to simulate social or demographic groups Hu and Collier (2024). This is commonly done with persona or sociodemographic (SD) prompting Hu and Collier (2024); Beck et al. (2024) by both technical and non-technical practitioners alike Beck et al. (2024).
Prior work on SD prompting has typically suffered from one or more key shortcomings. First, many studies have assumed a single majority answer for each sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025); Mukherjee et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. In addition, prompting for a distribution, rather than a representative response, could potentially make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2024); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024). Other work seeking to usefully replicate ground-truth answer distributions have been agnostic to underlying demographics, i.e., evaluating LLM distributions prompted with specific demographics with ground-truth distributions of multiple other demographics Beck et al. (2024).
In this work, we evaluate SD prompting in a distributional manner across three large survey datasets (both US and global), using a variety of methods to extract distributions from LLMs. We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs. We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups.
Concretely, we evaluate and explore distributional opinion alignment by studying the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data? Is this consistent across models, datasets, and distribution elicitation methods? (RQ 3) Can generated distributions be more easily aligned with some SD groups over others? What is the effect of calibration for individual groups? (RQ 4) Does calibration work with fewer supervised training examples? And (RQ 5) What are the effects of post-training (methods like SFT and DPO) on LLM-generated distributions?
Our experiments evaluate the alignment of LLM-generated distributions to human opinion, across three distribution elicitation methods, and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We find that while baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1), our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across settings (RQ2). Moreover, as few as 5 gold examples can be used to calibrate distributions (RQ4), though alignment to some demographics is degraded more than others (RQ3). In regard to LLM post-training (RQ5), the effect appears to be dataset dependent, though we find that calibrated alignment is less affected than uncalibrated distributions.
We conclude with concrete suggestions to practitioners seeking to predict diverse human responses (e.g., public opinion surveys) via LLMs, offering appropriate caution.
2 Related Work
Sociodemographic (SD) Prompting
To support safe deployment of models, it is important that modeling and evaluation practices account for social dimensions of diversity Santurkar et al. (2023) that shape our social identities, including sociodemographic factors such as age, gender, and ethnicity Hwang et al. (2023); Movva et al. (2024). SD prompting refers to incorporating demographic variables into input prompts to enhance model alignment with desired outputs Alipour et al. (2024); Joshi et al. (2025). As past work has shown that neither annotator agreement methodologies nor reward mechanisms often produce significant differences in the solution Yang et al. (2024), prompting methods may support more efficient processes. Some work has reported that SD prompting can be effective in improving model performance Beck et al. (2024), supporting SD-specific diversification of predictions without need for collecting sensitive annotator SD information Gupta et al. (2023).
Pitfalls of Sociodemographic Prompting
Prior studies evaluating sociodemographic prompting have reported mixed results in effectiveness in accurately simulating a social group Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024). While some works show that sociodemographic prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024).
Other work warns against such methods due to misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2024); Cheng et al. (2023b). Simply incorporating demographic features into prompts does not always enable LLMs to adopt the perspectives of specific demographic groups Sun et al. (2025). Furthermore, a vast body of work has established biases in language models that may be perpetuated with sociodemographic prompting Nadeem et al. (2020); Gallegos et al. (2024); Gupta et al. (2024). There is risk of "pigeonholing particular demographic groups into specific narratives" Cheng et al. (2023a).
Distribution Elicitation
Confidence elicitation is the process of estimating LLM’s confidence in their responses without model fine-tuning or accessing internal information Geng et al. (2024a). Recent research has explored methods for confidence elicitation which are more suitable for both closed-source commercial APIs as well as open-source models Lin et al. (2024). Similarly, distribution elicitation methods are techniques used to obtain probability distributions from experts about unknown quantities Falconer et al. (2022).
3 Approach
3.1 Datasets
We consider three human survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA), and the World Values Survey (WVS). Both OQA and WVS have been used to study sociodemographic prompting in prior work Santurkar et al. (2023); Durmus et al. (2024). These three datasets cover a diverse range of subjective topics, such as perceptions of science and public health, public opinion on gun control, data privacy, and various moral opinions and values. To reduce downstream LLM costs, we subset datasets to include all WGM ordinal questions and two questions per category from OQA and WVS. This totals 92 questions across WGM (14), OQA (38), and WVS (40).
Demographics for each dataset are shown in Table 11 (Appendix). Over all datasets, questions, and demographics, there are 4,500 human response distributions. For each question, model, and distribution elicitation method (Section 3.2), we predict the probability distribution of human responses over answer choices. This yields 220,500 generated SD-specific response distributions and 4,500 generated SD-agnostic response distributions.
3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To predict these distributions via LLM, we apply (with some modifications) methods inspired from confidence literature as follows. We investigate techniques that work most broadly, supporting use with both black-box and white-box LLMs. That said, Appendix A.3 reports a smaller scale study with log probabilities. Appendix B shows prompts used for all methods.
Verbalized. This is simply directly prompting the model to output a distribution in numbers Geng et al. (2024b); Tian et al. (2023). For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over the answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample times and average for variance reduction. If a model ever does not generate a distribution as requested, we simply discard that output. We also renormalize generated distributions that do not sum to 1.
Self-random. We prompt the model to output a single answer choice, sampling times Xiong et al. (2024), with a temperature of 0.7. We then create a distribution over the responses. For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice, such as 4 for a lot or 1 for not at all. Note that this method estimates the log probabilities for large Tian et al. (2023). In our experiments, we sample times.
Paraphrase. We also prompt the model to output a single answer choice but use different paraphrases of the prompt instead of sampling responses for the same prompt Xiong et al. (2024). We use paraphrases for each prompt.
3.3 Prompts
We use two categories of prompts, and for each category, we prompt with the different elicitation methods. See Appendix B for prompts used.
Base. Base prompts exclude any sociodemographic information, asking questions from the original dataset with formatting slightly changed for LLMs. See Appendix Table 12 for full prompts.
Sociodemographic (SD). These use a form similar to “Imagine you are {d}. Question: ”, where {d} is a demographic value. The exact prompt varies with demographic and elicitation method. See Appendix Table 13 for full prompts used.
3.4 Metric
We use the opinion alignment metric from Santurkar et al. (2023) to measure similarity between elicited distributions and ground truth distributions for all results. This metric makes use of 1 Wasserstein distance, or earth-mover’s distance. As noted by Santurkar et al. (2023), this metric takes into account the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges from 0 to 1, but in our results we show the value as a percentage.
3.5 Models
We evaluate 15 models ranging in openness (open-source, open-weight, black box), size, modality, and post-training. Model families include Claude, Llama, Mistral, OLMo-2, and Qwen. Our main results report on only the most powerful model in each family. Others are reported in Appendix A.1.
3.6 Calibration
We apply supervised regression to transform LLM-generated distributions. We split the LLM distributions into train, development, and test sets (60-20-20). Results in Section 6 show that far less supervision is typically needed in practice.
The input to regression is the LLM distribution and the output is the scaled distribution. We split the distributions by answer choice before normalizing across choices. More formally, for a regression model and LLM distribution parameterized by the values for each answer choice , we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice . I.e., we train on pairs of to learn transformed values on held-out test questions. We learn a regression model for each dataset-LLM-elicitation setting and do development set hyperparameter tuning and model selection to choose between linear regression (including lasso and ridge) and random forest regression.
| Model | Base prompt | Sociodemographic prompt | |||||||||||
| WGM | Claude-3.5-v2 | 66.2 | 85.2 | 59.5 | 85.1 | 89.3 | 87.7 | 65.4 | 84.4 | 61.4 | 84.2 | 89.0 | 89.3 |
| Llama-3.2-90B | 68.1 | 84.6 | 73.0 | 86.2 | 84.8 | 89.0 | 70.6 | 86.0 | 67.6 | 86.1 | 85.0 | 89.8 | |
| Mistral-large | 62.4 | 84.7 | 72.0 | 83.9 | 89.4 | 88.9 | 68.4 | 84.7 | 63.0 | 84.9 | 87.3 | 88.4 | |
| OLMo-2-7B-I | 59.6 | 81.5 | 67.7 | 80.4 | 59.8 | 85.1 | 62.3 | 82.5 | 64.5 | 82.7 | 69.8 | 84.6 | |
| Qwen-2.5-72B | 57.7 | 83.1 | 53.3 | 83.2 | 88.2 | 87.1 | 66.0 | 84.1 | 63.4 | 85.2 | 89.1 | 89.4 | |
| Average | 62.8 | 83.8 | 65.1 | 83.8 | 82.3 | 87.6 | 66.5 | 84.3 | 64.0 | 84.6 | 84.0 | 88.3 | |
| OQA | Claude-3.5-v2 | 70.5 | 88.8 | 72.5 | 90.4 | 91.7 | 91.7 | 76.1 | 89.9 | 73.3 | 89.6 | 91.9 | 91.6 |
| Llama-3.2-90B | 79.3 | 89.6 | 75.3 | 89.4 | 86.8 | 87.9 | 79.2 | 90.1 | 76.1 | 90.0 | 83.4 | 85.9 | |
| Mistral-large | 79.5 | 89.9 | 75.3 | 88.3 | 85.0 | 86.2 | 75.8 | 89.2 | 72.4 | 89.5 | 83.8 | 84.7 | |
| OLMo-2-7B-I | 72.6 | 89.0 | 72.3 | 88.4 | 65.4 | 79.9 | 72.8 | 88.4 | 70.5 | 88.6 | 68.5 | 81.5 | |
| Qwen-2.5-72B | 73.9 | 89.7 | 67.0 | 88.8 | 88.4 | 87.9 | 74.4 | 90.0 | 71.3 | 89.5 | 89.2 | 88.6 | |
| Average | 75.2 | 89.4 | 72.5 | 89.1 | 83.5 | 86.7 | 75.7 | 89.5 | 72.7 | 89.4 | 83.4 | 86.5 | |
| WVS | Claude-3.5-v2 | 46.5 | 80.3 | 51.0 | 80.3 | 75.2 | 80.3 | 61.0 | 80.4 | 56.8 | 80.4 | 75.6 | 81.7 |
| Llama-3.2-90B | 61.6 | 79.9 | 59.1 | 80.3 | 64.6 | 80.3 | 62.1 | 80.8 | 59.5 | 81.5 | 67.7 | 82.7 | |
| Mistral-large | 48.8 | 80.3 | 44.0 | 82.2 | 72.8 | 80.3 | 54.3 | 80.4 | 51.5 | 80.3 | 76.6 | 83.8 | |
| OLMo-2-7B-I | 75.3 | 80.3 | 74.9 | 80.3 | 86.6 | 86.5 | 58.3 | 79.2 | 60.0 | 77.2 | 86.0 | 89.8 | |
| Qwen-2.5-72B | 39.6 | 82.0 | 42.4 | 82.3 | 74.0 | 77.9 | 49.1 | 82.4 | 49.4 | 81.5 | 73.4 | 81.7 | |
| Average | 54.4 | 80.6 | 54.3 | 81.1 | 74.6 | 81.1 | 57.0 | 80.6 | 55.4 | 80.2 | 75.9 | 83.9 | |
| Average | Claude-3.5-v2 | 61.1 | 84.8 | 61.0 | 85.3 | 85.4 | 86.6 | 67.5 | 84.9 | 63.8 | 84.7 | 85.5 | 87.5 |
| Llama-3.2-90B | 69.7 | 84.7 | 69.1 | 85.3 | 78.7 | 85.7 | 70.6 | 85.6 | 67.7 | 85.9 | 78.7 | 86.1 | |
| Mistral-large | 63.6 | 85.0 | 63.8 | 84.8 | 82.4 | 85.1 | 66.2 | 84.8 | 62.3 | 84.9 | 82.6 | 85.6 | |
| OLMo-2-7B-I | 69.2 | 83.6 | 71.6 | 83.0 | 70.6 | 83.8 | 64.5 | 83.4 | 65.0 | 82.8 | 74.8 | 85.3 | |
| Qwen-2.5-72B | 57.1 | 84.9 | 54.2 | 84.8 | 83.5 | 84.3 | 63.2 | 85.5 | 61.4 | 85.4 | 83.9 | 86.6 | |
| Average | 64.1 | 84.6 | 64.0 | 84.6 | 80.1 | 85.1 | 66.4 | 84.8 | 64.0 | 84.7 | 81.1 | 86.2 | |
4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts, and calibrating LLM-generated distributions using some human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The full results for all models are shown in Appendix Table 5.
Findings (RQ1, RQ2):
(1) the effectiveness of baseline SD prompting in generating aligned distributions is model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al. (2024); (2) calibration increases opinion alignment in aggregate across models, datasets, and elicitation methods by an average of 16.3%; and (3) calibration reduces variance within and across settings.
4.1 Does SD prompting improve alignment?
To study RQ1, we compare SD vs. base prompts. What is the effect of adding SD information on opinion alignment? Note that while prior work studies this question only for majority-voted responses Beck et al. (2024); Hu and Collier (2024), we instead look at the effect of adding SD information when evaluating distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets. The results for all models are shown in Appendix Table 5.
For both base and SD prompts, LLM-generated distributions are generally best with verbalized elicitation (Table 1), though with exceptions (e.g., OLMo-Instruct on the OQA dataset). As shown in the table, prompting with SD information does not necessarily increase opinion alignment with human responses. In fact, it is often comparable or even lower than prompting without any SD information. The degree of effectiveness in creating distributions that align with human opinion appears to be model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al. (2024).
4.2 Can we calibrate LLM distributions?
We have seen that prompting LLMs with SD information does not consistently increase opinion alignment with human responses. Since LLM-generated distributions have fairly low opinion alignment, with RQ2, we question whether we can calibrate these distributions to better align them with human response distributions on these survey datasets.
The results after regression, compared to pre-regression results, are shown in Table 1, with results for all models in Appendix Table 5. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, and by an average of 16.3%. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment decreases by an average of 4.1%. These results suggest that LLM-generated distributions for predicting human responses are somewhat uncalibrated, and a simple supervised regression can lead to higher opinion alignment.
After calibration, opinion alignment in each setting also has much lower variance. Calibrated distributions tend to have a much lower standard deviation; standard deviation is lower in 87.2% of dataset-LLM-elicitation settings, and is on average 1.62 times lower. That lower standard deviation leads to higher opinion alignment provides more evidence that LLMs might be exaggerating differences in opinion distributions between demographic groups Cheng et al. (2023a, b), and that calibration could be helpful in mitigating this.
In addition, we find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. We see that the standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. While prior work Beck et al. (2024); Hu and Collier (2024) and our own previous results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly. This means that with calibration, the choice of any particular model becomes less important.
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| World region (WGM) | Aus/NZ | 82.3 | 79.7 | 94.7 | 83.5 | 88.2 | 86.6 | 73.5 | 79.1 | 89.1 | 84.2 | 85.6 | 82.6 |
| Central Africa | 85.5 | 87.7 | 71.2 | 81.2 | 78.1 | 78.6 | 38.0 | 79.0 | 75.6 | 79.2 | 69.7 | 81.1 | |
| Cent. America & Mex. | 89.2 | 90.0 | 74.8 | 85.1 | 81.7 | 82.8 | 46.4 | 84.7 | 78.5 | 81.8 | 74.1 | 84.9 | |
| Central Asia | 82.5 | 80.2 | 88.9 | 82.8 | 87.7 | 85.3 | 75.5 | 79.9 | 88.8 | 84.4 | 84.7 | 82.5 | |
| East Asia | 88.4 | 85.7 | 86.8 | 85.7 | 88.7 | 88.6 | 61.6 | 82.0 | 86.4 | 82.8 | 82.4 | 85.0 | |
| Eastern Africa | 88.2 | 88.4 | 80.0 | 87.7 | 88.0 | 87.3 | 63.3 | 87.7 | 86.0 | 86.2 | 81.1 | 87.5 | |
| Eastern Europe | 93.0 | 88.4 | 85.4 | 88.1 | 89.6 | 89.4 | 54.6 | 81.9 | 89.9 | 88.5 | 82.5 | 87.3 | |
| Middle East | 91.1 | 87.9 | 87.7 | 92.9 | 92.5 | 91.7 | 59.0 | 85.1 | 91.4 | 88.8 | 84.3 | 89.3 | |
| North Africa | 90.0 | 87.5 | 81.6 | 87.8 | 86.9 | 87.2 | 53.0 | 81.4 | 85.1 | 84.1 | 79.3 | 85.6 | |
| Northern America | 83.2 | 80.9 | 91.7 | 85.0 | 88.2 | 87.0 | 68.4 | 82.4 | 92.4 | 88.8 | 84.8 | 84.8 | |
| Northern Europe | 82.8 | 80.6 | 94.7 | 84.8 | 89.6 | 87.6 | 72.4 | 81.0 | 91.3 | 86.5 | 86.2 | 84.1 | |
| South America | 88.9 | 89.5 | 74.0 | 84.4 | 80.9 | 82.0 | 45.1 | 82.5 | 77.7 | 81.0 | 73.3 | 83.9 | |
| South Asia | 84.2 | 84.0 | 86.7 | 84.8 | 88.9 | 87.3 | 70.5 | 83.7 | 87.0 | 83.3 | 83.5 | 84.6 | |
| Southeast Asia | 82.1 | 82.5 | 79.1 | 80.2 | 86.3 | 84.6 | 70.6 | 81.9 | 84.2 | 81.7 | 80.5 | 82.2 | |
| Southern Africa | 86.4 | 90.1 | 74.2 | 84.4 | 82.4 | 82.8 | 49.1 | 88.8 | 79.7 | 83.3 | 74.4 | 85.9 | |
| Southern Europe | 93.2 | 89.1 | 84.5 | 88.7 | 90.0 | 89.7 | 53.4 | 81.7 | 88.5 | 88.2 | 81.9 | 87.5 | |
| Western Africa | 90.7 | 90.4 | 81.0 | 91.6 | 88.1 | 88.2 | 57.7 | 90.3 | 85.7 | 86.4 | 80.6 | 89.4 | |
| Western Europe | 83.2 | 80.6 | 93.0 | 86.4 | 90.3 | 88.5 | 71.6 | 84.2 | 93.5 | 88.7 | 86.3 | 85.7 | |
| All Demographics (WGM) | 89.3 | 87.7 | 84.8 | 89.0 | 89.4 | 88.9 | 59.8 | 85.1 | 88.2 | 87.1 | 82.3 | 87.5 | |
| P.I.(OQA) | Very conservative | 85.1 | 84.8 | 77.4 | 79.0 | 76.1 | 77.9 | 63.8 | 77.5 | 81.7 | 81.0 | 76.8 | 80.0 |
| Conservative | 89.0 | 89.0 | 82.4 | 83.7 | 81.3 | 82.9 | 64.4 | 79.2 | 85.7 | 85.1 | 80.6 | 84.0 | |
| Moderate | 92.3 | 92.3 | 87.9 | 88.9 | 85.7 | 87.0 | 65.0 | 79.5 | 88.5 | 88.0 | 83.9 | 87.1 | |
| Liberal | 91.6 | 92.1 | 89.5 | 89.3 | 85.5 | 85.9 | 63.1 | 77.8 | 88.6 | 88.9 | 83.7 | 86.8 | |
| Very liberal | 87.8 | 88.3 | 86.5 | 85.9 | 83.3 | 83.3 | 61.1 | 76.1 | 84.8 | 85.0 | 80.7 | 83.7 | |
| All Demographics (OQA) | 91.7 | 91.7 | 86.8 | 87.9 | 85.0 | 86.2 | 65.4 | 79.9 | 88.4 | 87.9 | 83.5 | 86.7 | |
| I.(WVS) | High | 77.1 | 82.2 | 66.0 | 82.2 | 74.3 | 82.2 | 87.6 | 89.9 | 75.9 | 79.8 | 76.2 | 83.3 |
| Middle | 75.2 | 80.2 | 64.7 | 80.2 | 72.7 | 80.2 | 86.8 | 86.6 | 73.6 | 77.8 | 74.6 | 81.0 | |
| Low | 77.4 | 82.4 | 66.8 | 82.4 | 75.0 | 82.4 | 87.8 | 88.5 | 76.1 | 80.0 | 76.6 | 83.1 | |
| All Demographics (WVS) | 75.2 | 80.3 | 64.6 | 80.3 | 72.8 | 80.3 | 86.6 | 86.5 | 73.9 | 77.9 | 74.6 | 81.1 | |
5 How does alignment vary across SDs?
So far, we have studied how well LLMs can predict responses from SD groups aggregated across all groups, with the goal of comparing different models and distribution elicitation methods. However, a particular model or method might be more aligned with some sociodemographics over others. With RQ3, we seek to understand the differences at the more granular demographic level.
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it performed the best with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. We choose 5 out of 15 LLMs; the most recent/powerful from each model family: Claude 3.5 v2, Llama 3.2 90B, Mistral large, OLMo-2 7B Instruct, and Qwen 2.5 72B. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 2. See Appendix A.4 for alignment for other demographics.
Findings (RQ3):
(1) Calibration produces distributions more aligned with some demographics over others; (2) Claude models are more highly aligned with OQA dataset SD groups; (3) different models are better aligned with specific demographics.
5.1 How does calibration affect alignment?
First, we compare opinion alignment between base and calibrated distributions. We use the regression models trained on all sociodemographic groups and study how such aggregate regression models might calibrate individual sociodemographics. We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more aligned for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. As expected, calibration increases alignment for some demographics and reduces alignment for others. We find that it tends to increase alignment of SD groups that were less aligned with base-prompted distributions, such as Central Africa, Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment is decreased for very high aligned demographic groups, such as Aus/NZ and Northern Europe with Llama 90B.
5.2 How does alignment vary across LLMs?
Next, we compare the opinion alignment between different LLMs. For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East. Interestingly, Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data. For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen. In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. This is discussed further in Section 7. All models have lowest alignment with middle income populations. This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income. In other words, income might be less of an important factor in determining the answer to the survey questions for those with middle income.
6 Does supervision work with less data?
Although we used 80% of our data as supervision for our regression models to calibrate the LLM-generated distributions, we study how opinion alignment varies with smaller amounts of supervised data to explore RQ4.
Findings (RQ4):
(1) As few as 5 training examples to the regression model can suffice to achieve close to minimal MSE; (2) on average, degradation in alignment for individual demographics is close to zero at 5 examples; (3) demographics with the largest opinion alignment degradation is dependent on base vs. SD prompts; and (4) regression models generalize fairly well to unseen datasets.
6.1 How much supervised data do you need?
We evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset. We average over 10 different random samples for each training data size. We find that as few as 5 examples (with 4 answer choices per example) can be enough to achieve close to the minimal MSE for any particular dataset-LLM-elicitation setting. We plot Mean Squared Error (MSE) over training data size in Appendix Figure 4 showing that MSE usually converges at 5 examples, though this is model and dataset dependent.
6.2 How are individual SD groups affected?
How does using a small set of random examples affect alignment of individual demographics? Although degradation on average is close to zero, individual demographics are affected differently. Table 3 shows those most affected and the degradation amount for both base and SD-prompted verbalized distributions. We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant. Looking at absolute values, using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
| Category | Demographic | ||||
| Base | Region | Southeast Asia | 81.0 | 79.4 | 1.6 |
| Education | Less than HS | 83.1 | 81.6 | 1.5 | |
| Race | Black | 86.6 | 85.1 | 1.5 | |
| Race | Hispanic | 87.2 | 85.9 | 1.4 | |
| Pol. Party | Democrat | 84.7 | 83.3 | 1.4 | |
| SD | Region | Northern Europe | 87.1 | 84.6 | 2.4 |
| Region | Western Europe | 87.1 | 84.9 | 2.3 | |
| Region | Aus/NZ | 84.3 | 82.1 | 2.2 | |
| Education | Tertiary | 88.6 | 86.5 | 2.1 | |
| Employment | Unemployed | 86.9 | 84.9 | 2.0 |
6.3 Do models generalize out-of-domain?
We also study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset. We train regression models on two of our three datasets and evaluate on the held-out dataset. Results are shown in Appendix A.5. As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset. Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM. This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
7 Does post-training impact alignment?
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + SFT, base + SFT + DPO, and the instruct model (base + SFT + DPO + RLVR). We study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it generally is best aligned.
Findings (RQ5):
Calibrated alignment is similar both before and after each phase of post-training, though pre-calibrated alignment decreases for 2/3 datasets when post-training is added.
7.1 What is the effect of calibration?
Figure 3 plots opinion alignment for all OLMo models. Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets. OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to affect opinion alignment much less.
7.2 What is the effect of post-training?
Changes in alignment after various post-training methods is dataset-dependent. For WGM, alignment of the base distribution falls significantly after post-training, though alignment for the SD distributions slightly increase with RLVR. Calibration alignment stays consistent after post-training. On OQA, alignment decreases with SFT, but stays about the same when adding other post-training. This is true of base, SD, and calibrated distributions, though calibration alignment decreases the least. On WVS however, alignment increases after post-training methods are added, and alignment changes less with calibration.
We suspect the increase in performance on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS. The drop in performance for uncalibrated distributions on WGM and OQA might indicate that OLMO’s post-training data is less aligned with some of the populations of these datasets.
8 Suggestions for practitioners
One must acknowledge risks and and exercise caution in using LLMs to approximate responses from diverse SD groups, since fallible LLMs may naturally exhibit random errors and systematic biases in their outputs. That said, the ease, speed, and cost of automated approximation may motivate some practitioners in some settings to use LLMs to approximate human responses, provided the predictive accuracy is “good enough” for their needs. Whether this is so will depend greatly on the varying accuracy needs of different real-world uses cases, as well as predictive accuracy relative to the SD groups of interest to each use case. We also note that such approximation may represent only an initial starting point, providing an initial guess of the data while awaiting for the real human response data to more slowly trickle in from participants.
In general, we suggest that practitioners: (1) evaluate distributions rather than a single response when evaluating alignment with human responses, and (2) calibrate LLM-generated distributions. As our results shown, calibration can be effective with as few as 5 supervised examples, though degradation varies different SD groups, so one must acknowledge the tradeoffs between risk, cost and quality in performing calibration with less supervision. We provide a summary of our results for the five models shown in Table 1.
8.1 Which methods and models work best?
As shown in Table 1, although the best distribution elicitation method is LLM and dataset dependent, verbalized elicitation leads to the best aligned distributions in the majority of settings and has a higher average across settings. Although this is true both pre-calibration and post-calibration, all methods are much more closely aligned post-calibration: without calibration, verbalized elicitation is the most aligned in 90% of settings, and with calibration it is the most aligned in 63% of settings. Adding sociodemographic information does not make LLM distributions more aligned with human distributions consistently; SD verbalized distributions are more aligned than their base verbalized counterparts in 67% of settings pre-calibration and 60% of settings post-calibration.
Overall, using verbalized elicitation (followed by calibration) might be best to obtain the most aligned distributions. However, we note that after calibration, the alignment of distributions produced by all the methods are fairly close (within 2% of each other). We also find that prompting with SD information is not as important, though might lead to slightly more aligned distributions in aggregate.
The most aligned models are much more dataset dependent. Claude 3.5 v2 is most aligned for WGM and OQA pre-calibration (with and without SD information). Post-calibration, Claude is most aligned for OQA, while Llama 3.2 90B is most aligned for WGM. For WVS, OLMo 7B Instruct is most aligned (both pre/post calibration).
9 Discussion
Essentializing sociodemographics. We note that by prompting LLMs with SD information – and evaluating their generated distributions against human participants with those sociodemographics– we are essentializing the sociodemographic to the identities of the survey respondents and their opinions Wang et al. (2024). However, practitioners still seek to understand opinions of specific populations (e.g., for election polling) where an aligned LLM might be beneficial for faster iteration on survey questions. Although sociodemographics are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023).
The effect of SD prompting. While our SD information provides valuable context about us, it alone does not determine our opinions, and it is fallacious to assume otherwise. As we have seen, prompting with SD information on its own also does not necessarily lead to more aligned distributions. Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al. (2023); Do et al. (2025). One could make personalized predictions for a sample of a group’s individuals members Gordon et al. (2022) rather than a single distributive prediction for the group. Multimodal prompts might also add SD information implicitly, though this might come with additional biases.
LLM distributions are uncalibrated. We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated – simple regression allows us to scale distributions to be much more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others. We also find that calibration reduces variance, both across examples and across settings. Since calibrated distributions generally have higher performance, this might provide evidence for LLMs exaggerating differences between sociodemographic groups; exaggerations that are reduced with calibration.
Calibration on individual SD groups. We do not find many significant differences between calibrated/uncalibrated distributions at the individual SD group level. Calibrating responses from each individual SD would likely improve alignment for that particular SD, which we do not do here. However, we caution efficacy may vary by demographic; some demographic traits might be more relevant to the response distributions of certain questions (e.g., urban vs. rural areas might affect gun rights questions more than marital status).
The dataset, LLM, and distribution elicitation method matter. As prior work has found when evaluating SD prompting on human majority responses Beck et al. (2024); Hu and Collier (2024), in this work we find that opinion alignment is dataset, LLM, and distribution elicitation method dependent. However, we find that verbalized elicitation tends to elicit probabilities that are most aligned with human distributions in the most settings, both before and after calibration. We find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is dataset-dependent.
10 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques. Our work also offers guidance for those using LLMs to predict human responses in practice, and our benchmark can enable and stimulate future research.
11 Limitations
Prompting methods. We do not study chain of thought or other prompting methods which might increase overall performance. We do not experiment with varied temperatures for our generations. For elicitation methods, though a larger might improve generated distributions, our choice of is reasonable given inference costs. We use for verbalized elicitation (as opposed to ) since we found little variance between runs. We also leave in-context learning or fine-tuning Orlikowski et al. (2025) with human responses to future work.
Logit-based elicitation methods. We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs. That said, we do also report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.3). We find that using log probabilities leads to the most aligned distributions in 10/24 or 42% of settings with calibration. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied. We only studied a subset of the demographics available in the survey datasets, and each demographic individually. Future work might study a larger subset of demographics, as well as intersectional demographics.
Regression. Our regression predicts each answer choice individually and then normalizes the predicted answers. We also optimize regression for MSE instead of the alignment metric. We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better. Future work might also ensemble multiple elicitation methods.
Ethics Statement
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for social good – for practitioners, NLP researchers, or social science researchers seeking to accurately approximate human responses for survey development, annotation tasks, or social science studies – this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
We also note that evaluating LLM responses against human responses of particular sociodemographic groups often assumes that everyone in a particular demographic group thinks the same way, and essentializes the demographic to an individual’s identity. In this work, we try to address this by evaluating against the distribution of human responses for a particular sociodemographic group. By doing so, we hope to better align LLMs with various sociodemographic groups in a distributionally pluralistic manner, rather than assume that all members of the group would answer the same way.
Acknowledgments
We thank Sooyong Lee for early contributions that did not make it into the ultimate paper. This research was supported in part by Cisco, Good Systems222https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), and a grant from Open Philanthropy. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix A Additional Results
A.1 Alignment results for all models
Results are in Table 5.
A.2 Minimal Supervision
Results are in Figure 4.
A.3 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024b) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 4.
| Model | Base prompt | SD prompt | |||
| WGM | Llama-3-70B | 73.1 | 86.2 | 67.5 | 86.9 |
| Llama-3.1-70B | 74.0 | 83.2 | 70.4 | 88.1 | |
| Llama-3.2-1B | 83.6 | 88.6 | 83.0 | 85.1 | |
| Llama-3.2-11B | 84.0 | 85.0 | 82.2 | 84.4 | |
| Average | 78.7 | 85.7 | 75.8 | 86.1 | |
| Std Dev | 5.9 | 2.3 | 8.0 | 1.7 | |
| OQA | Llama-3-70B | 75.4 | 85.2 | 72.3 | 88.1 |
| Llama-3.1-70B | 78.9 | 87.3 | 75.7 | 87.4 | |
| Llama-3.2-1B | 88.3 | 88.4 | 87.3 | 86.5 | |
| Llama-3.2-11B | 82.6 | 86.3 | 88.3 | 89.6 | |
| Average | 81.3 | 86.8 | 80.9 | 87.9 | |
| Std Dev | 5.5 | 1.4 | 8.1 | 1.3 | |
| WVS | Llama-3-70B | 54.1 | 82.8 | 51.1 | 83.5 |
| Llama-3.1-70B | 43.5 | 83.3 | 49.1 | 83.0 | |
| Llama-3.2-1B | 83.9 | 86.4 | 85.1 | 84.8 | |
| Llama-3.2-11B | 74.8 | 81.1 | 72.9 | 85.5 | |
| Average | 64.1 | 83.4 | 64.5 | 84.2 | |
| Std Dev | 18.6 | 2.2 | 17.4 | 1.1 | |
| Model | Base prompt | Sociodemographic prompt | |||||||||||
| WGM | OLMo-2-7B-Base | 65.6 | 80.1 | 79.2 | 81.8 | 82.2 | 83.9 | 74.8 | 81.9 | 71.0 | 81.8 | 74.9 | 82.9 |
| OLMo-2-7B-SFT | 80.6 | 81.8 | 70.3 | 81.8 | 68.6 | 86.3 | 75.6 | 82.6 | 70.2 | 82.1 | 65.4 | 85.0 | |
| OLMo-2-7B-DPO | 71.1 | 77.4 | 63.2 | 82.1 | 63.9 | 86.2 | 72.2 | 83.0 | 68.7 | 82.4 | 63.8 | 84.9 | |
| OLMo-2-7B-Instruct | 59.6 | 81.5 | 67.7 | 80.4 | 59.8 | 85.1 | 62.3 | 82.5 | 64.5 | 82.7 | 69.8 | 84.6 | |
| Llama-3-70B | 64.9 | 85.3 | 73.7 | 83.8 | 84.5 | 88.8 | 66.5 | 86.0 | 61.7 | 84.7 | 84.4 | 88.3 | |
| Llama-3.1-70B | 73.2 | 86.7 | 70.3 | 86.5 | 80.6 | 86.0 | 68.5 | 85.9 | 66.1 | 85.9 | 85.2 | 89.4 | |
| Llama-3.2-1B | 81.8 | 81.8 | 81.8 | 81.8 | – | – | 61.9 | 82.5 | 46.7 | 81.2 | 61.4 | 86.4 | |
| Llama-3.2-11B | 73.7 | 87.7 | 73.2 | 83.5 | 73.6 | 84.5 | 71.8 | 85.3 | 72.0 | 84.7 | 71.7 | 79.2 | |
| Llama-3.2-90B | 68.1 | 84.6 | 73.0 | 86.2 | 84.8 | 89.0 | 70.6 | 86.0 | 67.6 | 86.1 | 85.0 | 89.8 | |
| Qwen-2.5-72B | 57.7 | 83.1 | 53.3 | 83.2 | 88.2 | 87.1 | 66.0 | 84.1 | 63.4 | 85.2 | 89.1 | 89.4 | |
| Mistral-small | 61.6 | 83.9 | 61.6 | 84.6 | 88.9 | 88.8 | 67.5 | 86.6 | 64.4 | 85.8 | 89.0 | 89.1 | |
| Mistral-large | 62.4 | 84.7 | 72.0 | 83.9 | 89.4 | 88.9 | 68.4 | 84.7 | 63.0 | 84.9 | 87.3 | 88.4 | |
| Claude-3 | 57.8 | 81.1 | 50.0 | 81.8 | 84.8 | 86.7 | 64.9 | 80.6 | 52.9 | 81.7 | 88.4 | 88.1 | |
| Claude-3.5-v1 | 64.5 | 82.5 | 62.0 | 84.3 | 86.1 | 87.4 | 65.8 | 84.6 | 62.3 | 84.5 | 86.3 | 88.3 | |
| Claude-3.5-v2 | 66.2 | 85.2 | 59.5 | 85.1 | 89.3 | 87.7 | 65.4 | 84.4 | 61.4 | 84.2 | 89.0 | 89.3 | |
| Average | 67.3 | 83.2 | 67.4 | 83.4 | 80.3 | 86.9 | 68.1 | 84.0 | 63.7 | 83.9 | 79.4 | 86.9 | |
| Std Dev | 7.5 | 2.7 | 9.0 | 1.8 | 9.9 | 1.7 | 4.1 | 1.8 | 6.7 | 1.7 | 10.3 | 3.0 | |
| OQA | OLMo-2-7B-Base | 85.0 | 89.2 | 82.6 | 88.4 | 82.1 | 81.8 | 80.8 | 88.3 | 77.2 | 88.1 | 83.8 | 87.6 |
| OLMo-2-7B-SFT | 78.8 | 88.0 | 79.4 | 86.8 | 68.7 | 81.6 | 80.7 | 87.9 | 77.8 | 87.2 | 72.6 | 83.0 | |
| OLMo-2-7B-DPO | 79.3 | 88.6 | 81.5 | 86.7 | 69.3 | 82.8 | 82.1 | 88.4 | 78.7 | 87.7 | 72.3 | 82.7 | |
| OLMo-2-7B-Instruct | 72.6 | 89.0 | 72.3 | 88.4 | 65.4 | 79.9 | 72.8 | 88.4 | 70.5 | 88.6 | 68.5 | 81.5 | |
| Llama-3-70B | 72.2 | 87.5 | 76.8 | 89.1 | 81.8 | 85.0 | 74.4 | 88.9 | 70.1 | 88.9 | 79.5 | 83.7 | |
| Llama-3.1-70B | 73.5 | 88.8 | 70.5 | 87.7 | 83.8 | 86.4 | 76.4 | 89.5 | 71.7 | 89.4 | 83.6 | 86.6 | |
| Llama-3.2-1B | 82.3 | 88.4 | 88.4 | 88.4 | – | – | 72.6 | 89.0 | 70.2 | 88.6 | 83.6 | 84.9 | |
| Llama-3.2-11B | 84.3 | 89.4 | 75.5 | 86.4 | 65.8 | 77.2 | 75.7 | 87.2 | 74.8 | 87.0 | 83.0 | 86.9 | |
| Llama-3.2-90B | 79.3 | 89.6 | 75.3 | 89.4 | 86.8 | 87.9 | 79.2 | 90.1 | 76.1 | 90.0 | 83.4 | 85.9 | |
| Qwen-2.5-72B | 73.9 | 89.7 | 67.0 | 88.8 | 88.4 | 87.9 | 74.4 | 90.0 | 71.3 | 89.5 | 89.2 | 88.6 | |
| Mistral-small | 77.3 | 91.6 | 73.6 | 90.5 | 86.9 | 87.7 | 75.1 | 90.4 | 71.6 | 90.3 | 87.8 | 89.0 | |
| Mistral-large | 79.5 | 89.9 | 75.3 | 88.3 | 85.0 | 86.2 | 75.8 | 89.2 | 72.4 | 89.5 | 83.8 | 84.7 | |
| Claude-3 | 73.5 | 89.9 | 66.4 | 90.3 | 87.6 | 86.5 | 76.4 | 90.6 | 72.6 | 90.3 | 89.3 | 88.6 | |
| Claude-3.5-v1 | 70.5 | 86.9 | 69.6 | 88.6 | 89.4 | 89.0 | 76.1 | 90.3 | 72.8 | 89.8 | 91.0 | 90.8 | |
| Claude-3.5-v2 | 70.5 | 88.8 | 72.5 | 90.4 | 91.7 | 91.7 | 76.1 | 89.9 | 73.3 | 89.6 | 91.9 | 91.6 | |
| Average | 76.8 | 89.0 | 75.1 | 88.6 | 80.9 | 85.1 | 76.6 | 89.2 | 73.4 | 89.0 | 82.9 | 86.4 | |
| Std Dev | 4.8 | 1.1 | 6.0 | 1.3 | 9.4 | 4.0 | 2.9 | 1.0 | 2.8 | 1.1 | 7.0 | 3.0 | |
| WVS | OLMo-2-7B-Base | 74.5 | 80.3 | 70.7 | 80.1 | 84.5 | 80.3 | 78.2 | 81.4 | 78.2 | 81.5 | 80.8 | 81.9 |
| OLMo-2-7B-SFT | 69.7 | 79.2 | 68.1 | 79.1 | 87.2 | 87.2 | 72.9 | 80.6 | 73.1 | 80.7 | 84.9 | 89.0 | |
| OLMo-2-7B-DPO | 66.2 | 80.3 | 70.5 | 79.8 | 89.1 | 83.9 | 70.6 | 80.9 | 70.2 | 80.7 | 82.5 | 85.6 | |
| OLMo-2-7B-Instruct | 75.3 | 80.3 | 74.9 | 80.3 | 86.6 | 86.5 | 58.3 | 79.2 | 60.0 | 77.2 | 86.0 | 89.8 | |
| Llama-3-70B | 59.7 | 80.3 | 61.7 | 80.6 | 81.3 | 83.0 | 67.9 | 80.8 | 62.8 | 80.5 | 78.1 | 82.1 | |
| Llama-3.1-70B | 61.9 | 81.5 | 56.4 | 82.1 | 65.2 | 80.3 | 63.6 | 80.8 | 57.6 | 81.3 | 68.3 | 82.8 | |
| Llama-3.2-1B | 80.3 | 80.3 | 80.3 | 80.3 | – | – | 68.7 | 82.8 | 66.6 | 78.9 | – | – | |
| Llama-3.2-11B | 75.4 | 82.0 | 76.0 | 84.8 | 82.7 | 84.0 | 78.2 | 82.2 | 78.0 | 81.2 | 66.5 | 87.0 | |
| Llama-3.2-90B | 61.6 | 79.9 | 59.1 | 80.3 | 64.6 | 80.3 | 62.1 | 80.8 | 59.5 | 81.5 | 67.7 | 82.7 | |
| Qwen-2.5-72B | 39.6 | 82.0 | 42.4 | 82.3 | 74.0 | 77.9 | 49.1 | 82.4 | 49.4 | 81.5 | 73.4 | 81.7 | |
| Mistral-small | 42.9 | 80.3 | 46.3 | 80.3 | 75.0 | 80.3 | 48.0 | 80.4 | 46.4 | 80.4 | 68.6 | 81.7 | |
| Mistral-large | 48.8 | 80.3 | 44.0 | 82.2 | 72.8 | 80.3 | 54.3 | 80.4 | 51.5 | 80.3 | 76.6 | 83.8 | |
| Claude-3 | 47.3 | 80.3 | 55.7 | 80.3 | 74.1 | 80.3 | 55.3 | 80.2 | 57.3 | 79.8 | 73.2 | 81.7 | |
| Claude-3.5-v1 | 44.3 | 80.3 | 49.6 | 80.3 | 75.7 | 80.3 | 58.7 | 80.3 | 54.8 | 80.3 | 73.3 | 81.6 | |
| Claude-3.5-v2 | 46.5 | 80.3 | 51.0 | 80.3 | 75.2 | 80.3 | 61.0 | 80.4 | 56.8 | 80.4 | 75.6 | 81.7 | |
| Average | 59.6 | 80.5 | 60.4 | 80.9 | 77.7 | 81.8 | 63.1 | 80.9 | 61.5 | 80.4 | 75.4 | 83.8 | |
| Std Dev | 13.8 | 0.7 | 12.4 | 1.4 | 7.8 | 2.7 | 9.5 | 0.9 | 9.9 | 1.1 | 6.4 | 2.9 | |
A.4 Individual Sociodemographic Results
Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 6, 7, and 8 respectively.
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| Age | 15-29 | 95.2 | 92.7 | 85.0 | 94.1 | 92.9 | 92.5 | 56.6 | 88.6 | 90.2 | 90.7 | 84.0 | 91.7 |
| 30-49 | 92.7 | 90.7 | 85.7 | 93.1 | 93.1 | 92.7 | 58.9 | 87.0 | 90.6 | 89.8 | 84.2 | 90.7 | |
| 50+ | 88.6 | 87.5 | 87.9 | 91.0 | 91.4 | 90.7 | 63.7 | 85.9 | 90.6 | 87.1 | 84.4 | 88.4 | |
| Edu. | primary | 83.9 | 83.8 | 78.0 | 86.3 | 84.6 | 84.0 | 62.1 | 88.6 | 82.5 | 82.6 | 78.2 | 85.1 |
| secondary | 95.1 | 91.2 | 86.5 | 91.8 | 92.1 | 92.2 | 57.0 | 85.8 | 91.1 | 90.5 | 84.4 | 90.3 | |
| tertiary | 86.2 | 82.7 | 88.6 | 88.4 | 90.4 | 89.5 | 63.8 | 86.2 | 94.6 | 90.6 | 84.7 | 87.5 | |
| Employment | full time for employer | 92.3 | 88.5 | 89.1 | 91.1 | 93.4 | 92.7 | 60.6 | 86.0 | 93.9 | 90.8 | 85.9 | 89.8 |
| full time for self | 91.0 | 89.7 | 84.3 | 91.3 | 91.5 | 91.2 | 60.0 | 88.2 | 89.3 | 88.8 | 83.2 | 89.8 | |
| part time (no full time) | 91.2 | 89.9 | 86.0 | 93.2 | 92.4 | 92.1 | 60.9 | 88.1 | 90.2 | 89.2 | 84.1 | 90.5 | |
| part time (full time) | 92.1 | 90.6 | 83.1 | 93.0 | 90.3 | 90.4 | 58.5 | 89.7 | 88.0 | 87.9 | 82.4 | 90.3 | |
| out of work force | 91.1 | 89.6 | 85.6 | 92.1 | 91.5 | 91.1 | 60.5 | 86.7 | 89.7 | 88.6 | 83.7 | 89.6 | |
| unemployed | 92.9 | 92.7 | 80.3 | 91.4 | 88.1 | 88.8 | 51.6 | 87.4 | 85.0 | 87.5 | 79.6 | 89.6 | |
| Sex | female | 92.2 | 90.5 | 84.6 | 92.2 | 91.8 | 91.6 | 59.0 | 87.4 | 89.3 | 88.8 | 83.4 | 90.1 |
| male | 91.9 | 89.5 | 87.8 | 93.4 | 93.8 | 92.6 | 60.6 | 86.8 | 92.5 | 90.5 | 85.3 | 90.6 | |
| Income | fourth 20% | 92.9 | 90.7 | 87.5 | 93.3 | 94.3 | 93.2 | 60.1 | 86.9 | 92.0 | 90.3 | 85.4 | 90.9 |
| middle 20% | 91.9 | 90.3 | 86.2 | 92.5 | 92.7 | 92.4 | 60.2 | 87.1 | 90.7 | 89.4 | 84.3 | 90.3 | |
| poorest 20% | 90.2 | 88.8 | 82.5 | 91.3 | 89.2 | 89.1 | 59.0 | 87.9 | 86.9 | 87.2 | 81.6 | 88.9 | |
| second 20% | 91.0 | 89.5 | 85.0 | 91.7 | 91.4 | 91.1 | 60.4 | 87.3 | 89.4 | 88.4 | 83.4 | 89.6 | |
| top 20% | 92.5 | 89.1 | 87.5 | 92.5 | 93.6 | 92.7 | 59.5 | 86.8 | 93.7 | 92.0 | 85.4 | 90.6 | |
| Area | city/suburb | 94.2 | 91.4 | 85.9 | 92.7 | 93.0 | 93.2 | 56.4 | 86.3 | 90.4 | 90.3 | 84.0 | 90.8 |
| rural/small town | 90.3 | 89.2 | 86.2 | 92.0 | 92.0 | 91.5 | 62.1 | 87.8 | 90.2 | 88.5 | 84.2 | 89.8 | |
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| Age | 18-29 | 92.5 | 92.6 | 87.5 | 88.8 | 86.9 | 88.0 | 68.9 | 83.2 | 89.7 | 89.2 | 85.1 | 88.4 |
| 30-49 | 92.6 | 92.5 | 86.5 | 87.8 | 85.9 | 87.1 | 66.6 | 81.5 | 89.2 | 88.7 | 84.2 | 87.5 | |
| 50-64 | 92.5 | 92.4 | 86.5 | 87.8 | 84.1 | 85.8 | 63.5 | 78.6 | 88.1 | 87.5 | 82.9 | 86.4 | |
| 65+ | 92.1 | 92.1 | 87.3 | 88.3 | 83.9 | 85.4 | 62.1 | 76.9 | 87.9 | 87.3 | 82.7 | 86.0 | |
| Edu. | Associate’s degree | 92.5 | 92.4 | 85.9 | 87.4 | 84.6 | 86.5 | 65.3 | 80.5 | 88.3 | 87.6 | 83.3 | 86.9 |
| College graduate | 92.1 | 92.0 | 87.6 | 88.6 | 85.7 | 87.0 | 65.0 | 79.8 | 88.6 | 88.0 | 83.8 | 87.1 | |
| High school graduate | 91.5 | 91.4 | 85.2 | 86.5 | 83.2 | 84.8 | 63.6 | 78.5 | 87.8 | 87.2 | 82.3 | 85.7 | |
| Less than high school | 88.8 | 88.3 | 84.4 | 85.5 | 82.7 | 84.0 | 69.2 | 81.3 | 87.0 | 86.5 | 82.4 | 85.1 | |
| Postgraduate | 92.2 | 92.1 | 87.3 | 88.4 | 85.4 | 86.3 | 64.0 | 78.7 | 88.2 | 87.6 | 83.4 | 86.6 | |
| Some college, no degree | 93.7 | 93.7 | 87.0 | 88.4 | 84.9 | 86.5 | 65.2 | 80.2 | 89.4 | 88.8 | 84.0 | 87.5 | |
| Region | Midwest | 92.3 | 92.3 | 86.2 | 87.6 | 84.7 | 86.2 | 64.2 | 79.3 | 88.3 | 87.7 | 83.1 | 86.6 |
| Northeast | 92.9 | 92.8 | 88.2 | 89.1 | 85.6 | 87.1 | 64.0 | 78.7 | 89.1 | 88.6 | 84.0 | 87.3 | |
| South | 92.9 | 92.8 | 87.1 | 88.3 | 84.9 | 86.6 | 65.1 | 79.9 | 88.9 | 88.4 | 83.8 | 87.2 | |
| West | 92.8 | 92.7 | 87.9 | 89.1 | 86.0 | 87.3 | 65.0 | 80.1 | 89.1 | 88.5 | 84.2 | 87.5 | |
| Income | $100,000 or more | 90.7 | 90.6 | 85.7 | 86.9 | 84.1 | 85.3 | 64.0 | 78.6 | 86.9 | 86.2 | 82.3 | 85.5 |
| $30,000-$50,000 | 93.2 | 93.1 | 86.3 | 87.6 | 84.9 | 86.5 | 65.1 | 80.0 | 89.4 | 88.8 | 83.8 | 87.2 | |
| $50,000-$75,000 | 93.2 | 93.2 | 87.1 | 88.3 | 85.4 | 86.8 | 64.9 | 79.7 | 89.1 | 88.5 | 83.9 | 87.3 | |
| $75,000-$100,000 | 93.0 | 92.8 | 87.9 | 89.2 | 86.1 | 87.5 | 65.6 | 80.4 | 89.4 | 88.8 | 84.4 | 87.7 | |
| Less than $30,000 | 93.9 | 93.7 | 88.4 | 89.8 | 86.1 | 87.7 | 65.7 | 80.7 | 90.2 | 89.5 | 84.9 | 88.3 | |
| Marital | Divorced | 94.3 | 94.1 | 86.8 | 88.3 | 85.0 | 86.8 | 63.7 | 78.8 | 89.5 | 88.7 | 83.9 | 87.3 |
| Married | 92.0 | 91.9 | 86.5 | 87.8 | 84.7 | 86.1 | 64.6 | 79.5 | 88.3 | 87.7 | 83.2 | 86.6 | |
| Never been married | 93.3 | 93.2 | 88.5 | 89.9 | 87.5 | 88.5 | 66.5 | 81.4 | 89.9 | 89.3 | 85.1 | 88.5 | |
| Separated | 93.8 | 93.6 | 89.2 | 90.3 | 86.4 | 87.6 | 65.4 | 79.4 | 90.1 | 89.7 | 85.0 | 88.1 | |
| Widowed | 90.1 | 90.0 | 87.5 | 88.4 | 82.9 | 84.1 | 61.5 | 75.8 | 86.1 | 85.5 | 81.6 | 84.8 | |
| Pol. Party | Democrat | 90.5 | 91.0 | 89.5 | 89.2 | 85.3 | 85.8 | 62.5 | 77.3 | 87.9 | 88.4 | 83.1 | 86.3 |
| Independent | 92.1 | 92.0 | 86.3 | 87.4 | 84.3 | 85.7 | 64.6 | 79.6 | 88.2 | 87.6 | 83.1 | 86.5 | |
| Other | 91.2 | 91.0 | 84.5 | 86.1 | 84.0 | 85.6 | 65.8 | 80.9 | 87.5 | 86.7 | 82.6 | 86.1 | |
| Republican | 88.3 | 88.2 | 81.9 | 83.2 | 80.6 | 82.2 | 64.7 | 79.5 | 85.4 | 84.9 | 80.2 | 83.6 | |
| Race | Asian | 93.9 | 93.8 | 88.3 | 89.8 | 89.8 | 90.8 | 73.1 | 86.9 | 92.6 | 92.0 | 87.5 | 90.7 |
| Black | 93.4 | 93.6 | 91.4 | 91.8 | 87.4 | 88.1 | 65.7 | 80.3 | 90.5 | 90.0 | 85.7 | 88.8 | |
| Hispanic | 94.3 | 94.3 | 90.3 | 91.2 | 87.6 | 88.6 | 67.1 | 81.9 | 91.6 | 91.1 | 86.2 | 89.4 | |
| Other | 93.1 | 92.5 | 86.5 | 88.3 | 85.9 | 87.7 | 67.2 | 81.6 | 90.4 | 89.5 | 84.6 | 87.9 | |
| White | 92.2 | 92.1 | 86.4 | 87.7 | 84.5 | 85.9 | 63.8 | 78.8 | 88.2 | 87.6 | 83.0 | 86.4 | |
| Religion | Agnostic | 89.5 | 89.5 | 85.1 | 86.4 | 82.9 | 84.1 | 63.3 | 77.4 | 85.7 | 84.9 | 81.3 | 84.5 |
| Atheist | 87.6 | 88.1 | 85.3 | 85.9 | 82.4 | 83.2 | 62.4 | 76.8 | 84.6 | 84.1 | 80.5 | 83.6 | |
| Buddhist | 90.1 | 90.2 | 90.2 | 89.9 | 86.4 | 86.5 | 66.4 | 81.0 | 88.5 | 88.5 | 84.3 | 87.2 | |
| Hindu | 88.3 | 88.6 | 86.3 | 85.9 | 86.9 | 86.4 | 73.0 | 85.2 | 87.2 | 87.5 | 84.3 | 86.7 | |
| Jewish | 93.3 | 93.6 | 87.7 | 88.7 | 87.2 | 88.2 | 66.8 | 80.7 | 89.9 | 89.5 | 85.0 | 88.1 | |
| Mormon | 87.5 | 87.2 | 82.6 | 84.0 | 82.9 | 84.2 | 67.8 | 81.4 | 85.8 | 85.2 | 81.3 | 84.4 | |
| Muslim | 91.9 | 91.6 | 91.2 | 92.3 | 89.4 | 90.7 | 69.9 | 83.9 | 91.4 | 90.7 | 86.8 | 89.8 | |
| Nothing in particular | 93.2 | 93.1 | 89.2 | 90.3 | 86.6 | 88.0 | 65.1 | 80.2 | 89.4 | 89.0 | 84.7 | 88.1 | |
| Orthodox | 93.4 | 93.1 | 88.5 | 89.9 | 88.2 | 89.3 | 69.9 | 84.1 | 91.8 | 91.2 | 86.4 | 89.5 | |
| Other | 92.7 | 92.5 | 86.1 | 87.6 | 84.0 | 85.6 | 63.4 | 78.5 | 89.5 | 88.9 | 83.1 | 86.6 | |
| Protestant | 91.4 | 91.4 | 85.1 | 86.3 | 83.3 | 85.0 | 63.9 | 78.8 | 87.3 | 86.8 | 82.2 | 85.7 | |
| Roman Catholic | 93.1 | 93.0 | 87.5 | 88.7 | 85.7 | 87.2 | 66.3 | 80.8 | 89.3 | 88.7 | 84.4 | 87.7 | |
| Sex | Female | 93.4 | 93.3 | 87.8 | 89.1 | 85.1 | 86.5 | 63.5 | 78.4 | 89.2 | 88.7 | 83.8 | 87.2 |
| Male | 91.5 | 91.4 | 86.1 | 87.6 | 85.4 | 86.6 | 66.2 | 81.1 | 88.1 | 87.4 | 83.5 | 86.8 | |
| Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
| Age | 16-24 years | 76.4 | 81.6 | 65.5 | 81.6 | 73.9 | 81.6 | 86.2 | 88.3 | 75.3 | 79.2 | 75.5 | 82.5 |
| 25-34 years | 76.8 | 81.8 | 65.9 | 81.8 | 74.3 | 81.8 | 86.6 | 88.8 | 75.6 | 79.4 | 75.8 | 82.7 | |
| 35-44 years | 76.7 | 81.7 | 65.9 | 81.7 | 74.1 | 81.7 | 87.1 | 88.2 | 75.2 | 79.2 | 75.8 | 82.5 | |
| 45-54 years | 76.0 | 81.0 | 65.5 | 81.0 | 73.6 | 81.0 | 87.5 | 87.6 | 74.6 | 78.5 | 75.4 | 81.8 | |
| 55-64 years | 74.6 | 79.5 | 64.3 | 79.5 | 72.3 | 79.5 | 87.7 | 85.8 | 73.2 | 77.1 | 74.4 | 80.3 | |
| 65+ years | 72.9 | 77.9 | 62.8 | 77.9 | 70.8 | 77.9 | 87.5 | 84.6 | 71.6 | 75.5 | 73.1 | 78.8 | |
| Education | bachelor | 74.1 | 79.1 | 63.7 | 79.1 | 71.8 | 79.1 | 86.9 | 85.3 | 72.8 | 76.7 | 73.9 | 79.9 |
| doctoral | 73.3 | 78.8 | 62.6 | 78.8 | 71.0 | 78.8 | 86.0 | 85.7 | 72.5 | 76.4 | 73.1 | 79.7 | |
| early childhood | 78.7 | 83.7 | 68.1 | 83.7 | 76.4 | 83.7 | 90.2 | 91.3 | 77.4 | 81.2 | 78.2 | 84.7 | |
| lower secondary | 75.6 | 80.6 | 65.3 | 80.6 | 73.4 | 80.6 | 87.6 | 87.1 | 74.3 | 78.2 | 75.2 | 81.4 | |
| master | 74.3 | 79.5 | 63.6 | 79.5 | 72.0 | 79.5 | 86.8 | 87.0 | 73.2 | 77.1 | 74.0 | 80.5 | |
| post-secondary | 75.4 | 80.4 | 64.6 | 80.4 | 72.8 | 80.4 | 86.2 | 86.2 | 74.0 | 78.0 | 74.6 | 81.1 | |
| primary | 76.1 | 81.1 | 65.7 | 81.1 | 73.9 | 81.1 | 88.8 | 87.7 | 74.8 | 78.7 | 75.9 | 81.9 | |
| short-cycle tertiary | 73.7 | 78.7 | 63.2 | 78.7 | 71.4 | 78.7 | 85.9 | 85.2 | 72.3 | 76.2 | 73.3 | 79.5 | |
| upper secondary | 75.4 | 80.3 | 64.6 | 80.3 | 72.9 | 80.3 | 86.2 | 86.7 | 74.0 | 77.9 | 74.6 | 81.1 | |
| Employment | full time | 75.1 | 80.1 | 64.7 | 80.1 | 72.8 | 80.1 | 87.1 | 86.5 | 73.7 | 77.7 | 74.7 | 80.9 |
| housewife | 75.7 | 80.7 | 65.4 | 80.7 | 73.5 | 80.7 | 87.5 | 87.9 | 74.4 | 78.3 | 75.3 | 81.7 | |
| other | 74.8 | 79.8 | 64.7 | 79.8 | 72.7 | 79.8 | 87.1 | 83.9 | 73.6 | 77.4 | 74.6 | 80.1 | |
| part time | 76.8 | 81.8 | 66.2 | 81.8 | 74.3 | 81.8 | 87.1 | 88.2 | 75.3 | 79.4 | 75.9 | 82.6 | |
| retired/pensioned | 72.9 | 77.9 | 62.6 | 77.9 | 70.7 | 77.9 | 87.1 | 84.8 | 71.6 | 75.4 | 73.0 | 78.8 | |
| self-employed | 77.0 | 82.0 | 66.0 | 82.0 | 74.4 | 82.0 | 87.4 | 89.6 | 75.7 | 79.6 | 76.1 | 83.0 | |
| student | 75.6 | 80.8 | 64.9 | 80.8 | 73.1 | 80.8 | 86.1 | 87.0 | 74.4 | 78.4 | 74.8 | 81.6 | |
| unemployed | 78.0 | 83.0 | 67.0 | 83.0 | 75.4 | 83.0 | 86.9 | 89.0 | 76.8 | 80.6 | 76.8 | 83.7 | |
| Household size | 1 | 74.1 | 79.1 | 63.8 | 79.1 | 72.0 | 79.1 | 87.2 | 85.0 | 72.9 | 76.7 | 74.0 | 79.8 |
| 2 | 73.5 | 78.5 | 63.3 | 78.5 | 71.4 | 78.5 | 87.0 | 84.6 | 72.2 | 76.1 | 73.5 | 79.2 | |
| 3 | 75.0 | 80.0 | 64.5 | 80.0 | 72.6 | 80.0 | 87.1 | 85.9 | 73.4 | 77.5 | 74.5 | 80.7 | |
| 4 | 75.6 | 80.5 | 65.1 | 80.5 | 73.1 | 80.5 | 86.9 | 86.7 | 74.0 | 78.1 | 74.9 | 81.3 | |
| 5 | 76.5 | 81.4 | 65.5 | 81.4 | 73.9 | 81.4 | 87.0 | 88.8 | 75.2 | 79.0 | 75.6 | 82.4 | |
| 6 | 76.8 | 81.8 | 65.7 | 81.8 | 74.1 | 81.8 | 87.2 | 89.5 | 75.6 | 79.4 | 75.9 | 82.9 | |
| 7 persons or more | 78.9 | 86.5 | 66.1 | 86.5 | 74.6 | 86.5 | 80.6 | 86.0 | 78.4 | 84.8 | 75.7 | 86.1 | |
| Imm. | Immigrant | 71.7 | 76.7 | 61.3 | 76.7 | 69.5 | 76.7 | 86.0 | 83.8 | 70.4 | 74.2 | 71.8 | 77.6 |
| Native | 76.1 | 81.1 | 65.5 | 81.1 | 73.5 | 81.1 | 87.0 | 87.6 | 74.6 | 78.7 | 75.3 | 81.9 | |
| Marital | Divorced | 74.2 | 79.2 | 63.9 | 79.2 | 72.0 | 79.2 | 86.7 | 84.8 | 72.9 | 76.8 | 73.9 | 79.8 |
| Living together | 74.1 | 79.2 | 63.5 | 79.2 | 71.7 | 79.2 | 83.8 | 82.7 | 72.6 | 76.7 | 73.1 | 79.4 | |
| Married | 76.2 | 81.2 | 65.6 | 81.2 | 73.6 | 81.2 | 87.4 | 88.3 | 74.7 | 78.8 | 75.5 | 82.1 | |
| Separated | 74.5 | 79.4 | 63.9 | 79.4 | 72.1 | 79.4 | 84.9 | 83.6 | 73.1 | 77.0 | 73.7 | 79.8 | |
| Single | 75.9 | 81.0 | 65.3 | 81.0 | 73.5 | 81.0 | 86.6 | 87.1 | 74.5 | 78.6 | 75.2 | 81.7 | |
| Widowed | 75.0 | 80.0 | 64.2 | 80.0 | 72.5 | 80.0 | 87.0 | 86.5 | 73.7 | 77.5 | 74.5 | 80.8 | |
| Religion | Buddhist | 74.8 | 79.8 | 63.8 | 79.8 | 72.1 | 79.8 | 85.8 | 86.1 | 73.5 | 77.3 | 74.0 | 80.6 |
| Hindu | 79.7 | 86.0 | 68.4 | 86.0 | 76.9 | 86.0 | 85.3 | 93.4 | 79.6 | 83.6 | 78.0 | 87.0 | |
| Jew | 76.3 | 81.5 | 65.2 | 81.5 | 73.6 | 81.5 | 88.0 | 87.9 | 75.2 | 79.0 | 75.7 | 82.3 | |
| Muslim | 76.9 | 81.8 | 66.1 | 81.8 | 74.4 | 81.8 | 89.5 | 90.5 | 75.6 | 79.4 | 76.5 | 83.1 | |
| Orthodox | 73.1 | 78.6 | 62.3 | 78.6 | 70.7 | 78.6 | 85.5 | 85.8 | 72.4 | 76.2 | 72.8 | 79.6 | |
| Other | 70.8 | 75.7 | 61.4 | 75.7 | 68.3 | 75.7 | 85.7 | 80.1 | 69.2 | 73.3 | 71.1 | 76.1 | |
| Other Christian | 71.6 | 76.6 | 61.5 | 76.6 | 69.5 | 76.6 | 84.0 | 80.6 | 70.3 | 74.2 | 71.4 | 76.9 | |
| Protestant | 74.3 | 79.3 | 63.7 | 79.3 | 71.9 | 79.3 | 85.3 | 83.9 | 72.9 | 76.9 | 73.6 | 79.7 | |
| Roman Catholic | 73.9 | 78.9 | 63.4 | 78.9 | 71.6 | 78.9 | 85.2 | 83.8 | 72.6 | 76.5 | 73.3 | 79.4 | |
| none | 70.6 | 75.6 | 61.0 | 75.6 | 68.5 | 75.6 | 84.9 | 81.1 | 69.3 | 73.2 | 70.9 | 76.2 | |
| Sex | Female | 75.2 | 80.2 | 64.3 | 80.2 | 72.5 | 80.2 | 86.1 | 86.5 | 73.8 | 77.7 | 74.4 | 81.0 |
| Male | 76.3 | 81.2 | 65.9 | 81.2 | 73.9 | 81.2 | 87.8 | 87.7 | 74.8 | 78.8 | 75.7 | 82.0 | |
A.5 Generalization results
| Model | Base prompt | Sociodemographic prompt | |||||||||||
| WGM | OLMo-2-7B-Base | 65.6 | 82.4 | 79.2 | 81.8 | 82.2 | 82.6 | 74.8 | 81.9 | 71.0 | 82.1 | 74.9 | 82.6 |
| OLMo-2-7B-SFT | 80.6 | 81.8 | 70.3 | 81.3 | 68.6 | 83.9 | 75.6 | 82.4 | 70.2 | 82.2 | 65.4 | 82.7 | |
| OLMo-2-7B-DPO | 71.1 | 81.7 | 63.2 | 82.1 | 63.9 | 83.1 | 72.2 | 82.7 | 68.7 | 82.4 | 63.8 | 83.0 | |
| OLMo-2-7B-Instruct | 59.6 | 82.0 | 67.7 | 81.8 | 59.8 | 82.8 | 62.3 | 81.7 | 64.5 | 82.2 | 69.8 | 84.4 | |
| Llama-3-70B | 64.9 | 83.6 | 73.7 | 81.8 | 84.5 | 86.5 | 66.5 | 84.6 | 61.7 | 84.1 | 84.4 | 87.1 | |
| Llama-3.1-70B | 73.2 | 82.7 | 70.3 | 84.4 | 80.6 | 81.8 | 68.5 | 85.2 | 66.1 | 81.8 | 85.2 | 87.9 | |
| Llama-3.2-1B | 81.8 | 81.8 | 81.8 | 81.8 | – | – | 61.9 | 82.3 | 46.7 | 80.7 | 61.4 | 86.4 | |
| Llama-3.2-11B | 73.7 | 85.9 | 73.2 | 81.4 | 73.6 | 82.1 | 71.8 | 84.5 | 72.0 | 84.1 | 71.7 | 81.1 | |
| Llama-3.2-90B | 68.1 | 84.1 | 73.0 | 84.4 | 84.8 | 86.2 | 70.6 | 84.7 | 67.6 | 84.4 | 85.0 | 86.4 | |
| Qwen-2.5-72B | 57.7 | 83.5 | 53.3 | 83.0 | 88.2 | 86.7 | 66.0 | 81.8 | 63.4 | 83.7 | 89.1 | 85.9 | |
| Mistral-small | 61.6 | 84.1 | 61.6 | 83.7 | 88.9 | 85.7 | 67.5 | 84.6 | 64.4 | 83.9 | 89.0 | 85.0 | |
| Mistral-large | 62.4 | 84.4 | 72.0 | 83.5 | 89.4 | 86.8 | 68.4 | 84.6 | 63.0 | 84.2 | 87.3 | 86.2 | |
| Claude-3 | 57.8 | 82.2 | 50.0 | 81.8 | 84.8 | 84.3 | 64.9 | 82.2 | 52.9 | 81.6 | 88.4 | 85.6 | |
| Claude-3.5-v1 | 64.5 | 83.9 | 62.0 | 83.5 | 86.1 | 84.9 | 65.8 | 84.7 | 62.3 | 84.5 | 86.3 | 88.1 | |
| Claude-3.5-v2 | 66.2 | 84.3 | 59.5 | 83.1 | 89.3 | 86.5 | 65.4 | 84.3 | 61.4 | 84.1 | 89.0 | 87.7 | |
| Average | 67.3 | 83.2 | 67.4 | 82.6 | 80.3 | 84.6 | 68.1 | 83.5 | 63.7 | 83.1 | 79.4 | 85.3 | |
| Std Dev | 7.5 | 1.2 | 9.0 | 1.1 | 9.9 | 1.8 | 4.1 | 1.3 | 6.7 | 1.2 | 10.3 | 2.2 | |
| OQA | OLMo-2-7B-Base | 85.0 | 89.1 | 82.6 | 88.4 | 82.1 | 88.0 | 80.8 | 88.3 | 77.2 | 88.3 | 83.8 | 88.4 |
| OLMo-2-7B-SFT | 78.8 | 88.4 | 79.4 | 88.4 | 68.7 | 83.4 | 80.7 | 88.3 | 77.8 | 88.4 | 72.6 | 84.2 | |
| OLMo-2-7B-DPO | 79.3 | 88.4 | 81.5 | 88.3 | 69.3 | 83.6 | 82.1 | 88.2 | 78.7 | 88.3 | 72.3 | 85.4 | |
| OLMo-2-7B-Instruct | 72.6 | 88.6 | 72.3 | 88.4 | 65.4 | 82.5 | 72.8 | 88.5 | 70.5 | 88.6 | 68.5 | 81.4 | |
| Llama-3-70B | 72.2 | 88.4 | 76.8 | 88.4 | 81.8 | 88.4 | 74.4 | 88.7 | 70.1 | 89.0 | 79.5 | 88.8 | |
| Llama-3.1-70B | 73.5 | 88.9 | 70.5 | 88.9 | 83.8 | 88.4 | 76.4 | 89.1 | 71.7 | 90.4 | 83.6 | 89.6 | |
| Llama-3.2-1B | 82.3 | 88.4 | 88.4 | 88.4 | – | – | 72.6 | 89.1 | 70.2 | 88.5 | 83.6 | 87.6 | |
| Llama-3.2-11B | 84.3 | 88.4 | 75.5 | 87.7 | 65.8 | 88.2 | 75.7 | 87.8 | 74.8 | 87.7 | 83.0 | 88.7 | |
| Llama-3.2-90B | 79.3 | 89.4 | 75.3 | 88.4 | 86.8 | 88.4 | 79.2 | 88.9 | 76.1 | 89.0 | 83.4 | 88.8 | |
| Qwen-2.5-72B | 73.9 | 88.4 | 67.0 | 88.5 | 88.4 | 88.9 | 74.4 | 88.4 | 71.3 | 88.4 | 89.2 | 89.1 | |
| Mistral-small | 77.3 | 88.4 | 73.6 | 88.4 | 86.9 | 88.4 | 75.1 | 88.4 | 71.6 | 88.4 | 87.8 | 89.5 | |
| Mistral-large | 79.5 | 88.8 | 75.3 | 88.4 | 85.0 | 88.8 | 75.8 | 88.9 | 72.4 | 88.4 | 83.8 | 87.5 | |
| Claude-3 | 73.5 | 88.4 | 66.4 | 88.4 | 87.6 | 88.9 | 76.4 | 88.4 | 72.6 | 88.5 | 89.3 | 89.3 | |
| Claude-3.5-v1 | 70.5 | 88.4 | 69.6 | 88.4 | 89.4 | 88.4 | 76.1 | 89.5 | 72.8 | 89.7 | 91.0 | 91.5 | |
| Claude-3.5-v2 | 70.5 | 88.4 | 72.5 | 88.4 | 91.7 | 89.2 | 76.1 | 89.2 | 73.3 | 89.4 | 91.9 | 90.8 | |
| Average | 76.8 | 88.6 | 75.1 | 88.4 | 80.9 | 87.4 | 76.6 | 88.6 | 73.4 | 88.7 | 82.9 | 88.0 | |
| Std Dev | 4.8 | 0.3 | 6.0 | 0.2 | 9.4 | 2.3 | 2.9 | 0.5 | 2.8 | 0.7 | 7.0 | 2.6 | |
| WVS | OLMo-2-7B-Base | 74.5 | 81.2 | 70.7 | 80.3 | 84.5 | 84.4 | 78.2 | 81.1 | 78.2 | 81.0 | 80.8 | 83.0 |
| OLMo-2-7B-SFT | 69.7 | 79.1 | 68.1 | 79.2 | 87.2 | 85.6 | 72.9 | 80.3 | 73.1 | 80.8 | 84.9 | 87.0 | |
| OLMo-2-7B-DPO | 66.2 | 78.3 | 70.5 | 79.7 | 89.1 | 82.8 | 70.6 | 80.3 | 70.2 | 80.8 | 82.5 | 84.3 | |
| OLMo-2-7B-Instruct | 75.3 | 80.9 | 74.9 | 80.5 | 86.6 | 87.6 | 58.3 | 78.1 | 60.0 | 76.9 | 86.0 | 88.6 | |
| Llama-3-70B | 59.7 | 76.3 | 61.7 | 79.3 | 81.3 | 82.9 | 67.9 | 79.2 | 62.8 | 79.6 | 78.1 | 79.4 | |
| Llama-3.1-70B | 61.9 | 76.8 | 56.4 | 75.8 | 65.2 | 70.7 | 63.6 | 78.0 | 57.6 | 76.7 | 68.3 | 72.1 | |
| Llama-3.2-1B | 80.3 | 80.3 | 80.3 | 80.3 | – | – | 68.7 | 82.5 | 66.6 | 79.0 | – | – | |
| Llama-3.2-11B | 75.4 | 81.9 | 76.0 | 84.4 | 82.7 | 83.6 | 78.2 | 83.2 | 78.0 | 81.9 | 66.5 | 77.0 | |
| Llama-3.2-90B | 61.6 | 76.8 | 59.1 | 75.4 | 64.6 | 70.2 | 62.1 | 77.7 | 59.5 | 76.5 | 67.7 | 71.5 | |
| Qwen-2.5-72B | 39.6 | 71.7 | 42.4 | 74.3 | 74.0 | 74.3 | 49.1 | 71.8 | 49.4 | 74.7 | 73.4 | 73.8 | |
| Mistral-small | 42.9 | 73.0 | 46.3 | 74.4 | 75.0 | 75.8 | 48.0 | 72.5 | 46.4 | 74.6 | 68.6 | 71.0 | |
| Mistral-large | 48.8 | 74.2 | 44.0 | 76.4 | 72.8 | 74.8 | 54.3 | 74.4 | 51.5 | 76.4 | 76.6 | 78.3 | |
| Claude-3 | 47.3 | 71.4 | 55.7 | 76.3 | 74.1 | 74.0 | 55.3 | 74.8 | 57.3 | 77.1 | 73.2 | 73.3 | |
| Claude-3.5-v1 | 44.3 | 71.2 | 49.6 | 74.6 | 75.7 | 77.7 | 58.7 | 76.7 | 54.8 | 78.1 | 73.3 | 75.5 | |
| Claude-3.5-v2 | 46.5 | 72.1 | 51.0 | 75.5 | 75.2 | 75.2 | 61.0 | 78.2 | 56.8 | 78.4 | 75.6 | 75.6 | |
| Average | 59.6 | 76.3 | 60.4 | 77.8 | 77.7 | 78.5 | 63.1 | 77.9 | 61.5 | 78.2 | 75.4 | 77.9 | |
| Std Dev | 13.8 | 3.9 | 12.4 | 3.0 | 7.7 | 5.8 | 9.5 | 3.4 | 9.9 | 2.3 | 6.4 | 5.8 | |
| Model | Base prompt | SD prompt | |||
| WGM | Llama-3-70B | 73.1 | 84.6 | 67.5 | 86.1 |
| Llama-3.1-70B | 74.0 | 81.8 | 70.4 | 86.1 | |
| Llama-3.2-1B | 83.6 | 83.5 | 83.0 | 83.1 | |
| Llama-3.2-11B | 84.0 | 84.3 | 82.2 | 81.8 | |
| Average | 78.7 | 83.5 | 75.8 | 84.3 | |
| Std Dev | 5.9 | 1.3 | 8.0 | 2.2 | |
| OQA | Llama-3-70B | 75.4 | 88.4 | 72.3 | 88.5 |
| Llama-3.1-70B | 78.9 | 88.4 | 75.7 | 88.4 | |
| Llama-3.2-1B | 88.3 | 85.2 | 87.3 | 86.6 | |
| Llama-3.2-11B | 82.6 | 88.3 | 88.3 | 88.4 | |
| Average | 81.3 | 87.6 | 80.9 | 88.0 | |
| Std Dev | 5.5 | 1.6 | 8.1 | 0.9 | |
| WVS | Llama-3-70B | 54.1 | 74.8 | 51.1 | 72.1 |
| Llama-3.1-70B | 43.5 | 71.3 | 49.1 | 70.8 | |
| Llama-3.2-1B | 83.9 | 82.1 | 85.1 | 84.6 | |
| Llama-3.2-11B | 74.8 | 78.3 | 72.9 | 75.9 | |
| Average | 64.1 | 76.6 | 64.5 | 75.8 | |
| Std Dev | 18.5 | 4.6 | 17.4 | 6.2 | |
Appendix B Prompts (Reproducibility)
B.1 Demographics used
Demographics are shown in Table 11.
| Dataset | Example Questions | Demographics |
| WGM | How much do you, personally, know about science? Do you know a lot, some, not much, or nothing at all? | age, education, employment status, gender, income quintile, living area (city or rural), world region |
| How much do you trust doctors and nurses in your country? | ||
| OQA | How important, if at all, is being a gun owner to your overall identity? | age, education, income, marital status, political ideology, political party, race, US region, religion, sex |
| How much confidence, if any, do you have in elected officials to act in the best interests of the public? | ||
| WVS | Do you agree strongly, agree, disagree, or disagree strongly with the following statement? One of my main goals in life has been to make my parents proud. | age, education, employment status, household size, immigrant status, income, marital status, religion, sex |
| Do you agree, disagree or neither agree nor disagree with the following statement? When jobs are scarce, employers should give priority to people of this country over immigrants. |
B.2 Base prompts
Base prompts are shown in Table 12.
| Prompt type | Prompt template |
| Verbalized | What percentage of people would respond with {answer choices} to the following question? Question: ‘{q}’ Respond with a list of {l} numbers which represent the percentages for {answer choices} in that order. Expected response format: [, , , ,] Response: |
| Log probabilities | Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: {answer} |
| Self-random | Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 1 | Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 2 | How might someone respond to the following question? Question: ‘{q}’ Possible responses are {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 3 | Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 4 | Below you will be asked to answer some questions. Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 5 | Answer the following question. Question: ‘{q}’ Response options: {answer choices}. Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
B.3 Sociodemographic prompts
Sociodemographic prompts are shown in Table 13.
| Prompt type | Prompt template |
| Verbalized | What percentage of people whose {att} is {c} would respond with {answer choices} to the following question? Question: ’{q}’ Respond with a list of {l} numbers which represent the percentages for {answer choices} in that order. Expected response format: [, , , ,] Response: |
| Log probabilities | Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: {answer} |
| Self-random | Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 1 | Imagine your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 2 | How would a person whose {att} is {c} respond to the following question? Question: ‘{q}’ Possible responses are {answer choices } Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 3 | Person: {c} {att} Question: ‘{q}’ Response options: {answer choices } Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 4 | Below you will be asked to provide a short description of your {att} and then answer some questions. Description: I am {c} Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |
| Paraphrase 5 | Answer the following question as if your {att} is {c}. Would you respond with {answer choices} to the following question? Question: ‘{q}’ Respond with only a single integer number in the inclusive range [1,{l}], representing the index of your selected answer choice in {answer choices}. Response: |

### v2

@@
-Beyond Sociodemographic Prompting: Using Supervision to Align LLMs with Human Response Distributions
+Improving the Distributional Alignment of LLMs using Supervision
 Abstract
-The ability to accurately predict how different population groups would answer subjective questions would have great value.
-In this work, we show that use of relatively simple supervision can greatly improve language model alignment with diverse population groups, as measured over three datasets spanning various topics.
-Beyond evaluating average performance, we also report how alignment varies across specific groups.
-The simplicity and generality of our approach promotes easy adoption, while our broad findings provide useful guidance for when to use or not use our approach in practice.
-By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a useful benchmark to stimulate future research 111Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment.
+The ability to accurately align LLMs with human population groups on subjective questions would have great value.
+In this work, we show that use of simple supervision can greatly improve language model alignment with diverse population groups more consistently, as measured over three datasets spanning various topics.
+Beyond evaluating average alignment, we also report how alignment varies across specific groups.
+Our broad findings provide insights into the distributional alignment of LLMs with diverse population groups.
+By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a benchmark to stimulate future research.111Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment
 1 Introduction
-Human surveys elicit valuable public opinion on a wide range of topics, such as proposed legislation, candidates for office (i.e., election polls), marketing or outreach campaigns, commercial products and services, etc.
-Hayati et al.
-(2024); Santurkar et al.
+As LLMs are increasingly incorporated into real-world systems, it is more and more important to evaluate their alignment with human values and opinions.
+A growing body of work seeks to align LLMs with particular human population groups, with the goal of simulating their opinions or behaviors, by steering LLM generations towards those groups Aher et al.
+(2023); Sun et al.
+(2024); Dillion et al.
 (2023).
-Subjective annotation tasks in NLP also seek diverse, human judgments Sap et al.
-(2022); Biester et al.
-(2022); Pei and Jurgens (2023).
-Given the frequency, breadth, and importance of obtaining opinion data from diverse populations, there has been great interest in developing effective ways to conduct such surveys accurately and efficiently.
-Recent work has sought to simulate human responses using LLMs in many areas, such as in psychology, sociology, and economic studies Aher et al.
-(2023), NLP annotation Nasution and Onan (2024), and large-scale survey creation and testing Rothschild et al.
+One common method to do so is persona or sociodemographic (SD) prompting, which refers to defining a persona with demographic or description variables within a prompt, with the goal of modeling the behavior of that group using an LLM Hu and Collier (2024); Beck et al.
 (2024).
-More specifically, LLMs are being increasingly used to simulate social or demographic groups Hu and Collier (2024).
-This is commonly done with persona or sociodemographic (SD) prompting Hu and Collier (2024); Beck et al.
-(2024) by both technical and non-technical practitioners alike Beck et al.
+Recent work in NLP has sought to evaluate SD prompting by studying its performance on large surveys and polls, which contain many subjective questions, along with hundreds of thousands of human responses Santurkar et al.
+(2023); Durmus et al.
+(2024); Masoud et al.
+(2025); AlKhamissi et al.
 (2024).
-Prior work on SD prompting has typically suffered from one or more key shortcomings.
-First, many studies have assumed a single majority answer for each sociodemographic group in modeling and/or evaluation Hwang et al.
+However, evaluation of SD prompting has typically suffered from one or more key shortcomings.
+First, many studies have assumed a single majority answer to each question for each sociodemographic group in modeling and/or evaluation Hwang et al.
 (2023); Hu and Collier (2024); Sun et al.
-(2025); Mukherjee et al.
+(2025a); Mukherjee et al.
+(2024); Durmus et al.
+(2024); Masoud et al.
+(2025); AlKhamissi et al.
 (2024).
 The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement.
-In addition, prompting for a distribution, rather than a representative response, could potentially make a model less susceptible to caricature and stereotyping Cheng et al.
+Prompting for a distribution, rather than a representative response, could potentially also make a model less susceptible to caricature and stereotyping Cheng et al.
 (2023a); Wang et al.
-(2024); Cheng et al.
+(2025a); Cheng et al.
 (2023b), better aligning with the goals of pluralistic alignment Sorensen et al.
 (2024).
-Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al.
+However, Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al.
 (2023); Sun et al.
+(2024), or evaluate LLM distributions prompted for specific demographics in a demographic-agnostic manner Beck et al.
 (2024).
-Other work seeking to usefully replicate ground-truth answer distributions have been agnostic to underlying demographics, i.e., evaluating LLM distributions prompted with specific demographics with ground-truth distributions of multiple other demographics Beck et al.
-(2024).
-In this work, we evaluate SD prompting in a distributional manner across three large survey datasets (both US and global), using a variety of methods to extract distributions from LLMs.
-We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs.
+In this work, we evaluate SD prompting in a distributional manner across three survey datasets using a variety of methods to extract distributions from LLMs.
+We focus on methods that work most broadly, supporting use with open-source, open-model, and API-access LLMs.
 We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups.
-Concretely, we evaluate and explore distributional opinion alignment by studying the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion?
+We use supervised calibration to mean learning a transformation to a distribution to be more similar to human response distributions, using human responses as supervised labels (described in detail in Section 3.6).
+We use alignment to mean the degree to which opinions align between humans and LLMs (see Section 3.4).
+Concretely, we investigate distributional opinion alignment via the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion?
 (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data?
@@
 (RQ 4) Does calibration work with fewer supervised training examples?
-And (RQ 5) What are the effects of post-training (methods like SFT and DPO) on LLM-generated distributions?
-Our experiments evaluate the alignment of LLM-generated distributions to human opinion, across three distribution elicitation methods, and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods.
-We find that while baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1), our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across settings (RQ2).
-Moreover, as few as 5 gold examples can be used to calibrate distributions (RQ4), though alignment to some demographics is degraded more than others (RQ3).
-In regard to LLM post-training (RQ5), the effect appears to be dataset dependent, though we find that calibrated alignment is less affected than uncalibrated distributions.
-We conclude with concrete suggestions to practitioners seeking to predict diverse human responses (e.g., public opinion surveys) via LLMs, offering appropriate caution.
+And (RQ 5) What are the effects of post-training (doing fine-tuning and reinforcement learning) on LLM-generated distributions?
+We evaluate the alignment of LLM-generated distributions to human opinion across 3 distribution elicitation methods and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods.
+We first show that baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1).
+In contrast, our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across prompt methods, elicitation methods, datasets, and LLMs (RQ2), though it improves for some SD groups and degrades for others (RQ3).
+Moreover, as few as 1-10 gold examples can be used to calibrate distributions (RQ4).
+In regard to LLM post-training (RQ5), calibration results in more consistent behavior across various post-training methods, both with and without SD prompting; without calibration, alignment is much more dependent on both the dataset and the post-training method used.
+We conclude with a discussion of the methods and models that generate the distributions most aligned with human groups on subjective questions.
+Caution is warranted given varying success in aligning to different groups as well as recognition that such efforts risk essentializing sociodemographic categories if not applied with care.
 2 Related Work
-Sociodemographic (SD) Prompting
-To support safe deployment of models, it is important that modeling and evaluation practices account for social dimensions of diversity Santurkar et al.
-(2023) that shape our social identities, including sociodemographic factors such as age, gender, and ethnicity Hwang et al.
+Social alignment of LLMs
+LLMs are increasingly used to simulate social human groups in a variety of settings, such as psychology, sociology, and economic studies Aher et al.
+(2023); Dillion et al.
+(2023), survey development Rothschild et al.
+(2024); Jansen et al.
+(2023), NLP annotation Nasution and Onan (2024); Hashemi et al.
 (2024); Movva et al.
+(2024), and multi-agent interactive simulation Hou et al.
+(2025); Zhou et al.
 (2024).
-SD prompting refers to incorporating demographic variables into input prompts to enhance model alignment with desired outputs Alipour et al.
-(2024); Joshi et al.
+Many methods have been proposed to better align LLMs with human values and opinions Ma et al.
 (2024).
-As past work has shown that neither annotator agreement methodologies nor reward mechanisms often produce significant differences in the solution Yang et al.
-(2024), prompting methods may support more efficient processes.
-Some work has reported that SD prompting can be effective in improving model performance Beck et al.
-(2024), supporting SD-specific diversification of predictions without need for collecting sensitive annotator SD information Gupta et al.
+For example, persona or sociodemographic (SD) prompting defines a persona with demographic or description variables within a prompt to simulate or model groups Hu and Collier (2024); Beck et al.
 (2024).
-Pitfalls of Sociodemographic Prompting
-Prior studies evaluating sociodemographic prompting have reported mixed results in effectiveness in accurately simulating a social group Beck et al.
+Other methods include: using knowledge of a user’s previous opinions to model their future opinions Hwang et al.
+(2023); Gordon et al.
+(2022), using psychological scaffolds to improve model rationalization Joshi et al.
+(2025), self-alignment via the simulation of social scenes Pang et al.
+(2024), extracting persona vectors to mitigate certain character traits Chen et al.
+(2025), creating more descriptive value profiles Sorensen et al.
+(2025), using in-context learning Choenni and Shutova (2025), and fine-tuning LLMs followed by collection of a small amount of human data Krsteski et al.
+(2025).
+Evaluation frameworks have also been proposed for assessing alignment and social simulation Kang et al.
+(2025); Shi et al.
+(2025); Lutz et al.
+(2025).
+Pitfalls of social alignment
+Various limitations have been reported, such as LLMs being unreliable as personalized annotators Dong et al.
+(2024), unable to use different languages to steer generation towards groups Kwok et al.
+(2024), biased towards some groups Sun et al.
+(2025b), lacking stability and steerability under prompt variations Khan et al.
+(2025), and lacking the reasoning depth to simulate humans in social experiments Gao et al.
+(2025).
+Prior work evaluating SD prompting has also reported mixed results in accurately simulating social groups Beck et al.
 (2024); Santurkar et al.
-(2023); Hu and Collier (2024).
-While some works show that sociodemographic prompting can improve performance, results vary greatly by prompt, model, and task Beck et al.
+(2023); Hu and Collier (2024); Kaiser et al.
+(2025).
+While some works show that SD prompting can improve performance, results vary greatly by prompt, model, and task Beck et al.
 (2024); Mukherjee et al.
+(2024); Bisbee et al.
 (2024).
@@
 (2023a); Wang et al.
-(2024); Cheng et al.
-(2023b).
-Simply incorporating demographic features into prompts does not always enable LLMs to adopt the perspectives of specific demographic groups Sun et al.
+(2025a); Cheng et al.
+(2023b); Kitadai et al.
 (2024).
-Furthermore, a vast body of work has established biases in language models that may be perpetuated with sociodemographic prompting Nadeem et al.
+Simply incorporating demographic features into prompts does not always enable LLMs to adopt the desired perspectives of specified demographic groups Sun et al.
+(2025a).
+Furthermore, a vast body of work has established biases in language models that may be perpetuated with SD prompting Nadeem et al.
 (2021); Gallegos et al.
@@
 (2024).
-There is risk of "pigeonholing particular demographic groups into specific narratives" Cheng et al.
-(2023a).
-Distribution Elicitation
-Confidence elicitation is the process of estimating LLM’s confidence in their responses without model fine-tuning or accessing internal information Geng et al.
-(2024a).
-Recent research has explored methods for confidence elicitation which are more suitable for both closed-source commercial APIs as well as open-source models Lin et al.
+Probabilistic reasoning with LLMs
+Prior work in NLP has studied the use of LLMs for eliciting and working with uncertainty and probability distributions.
+Paruchuri et al.
+(2024) study whether LLMs can reason about probability distributions, finding that LLMs are able to answer questions about distributions, though performance varies based on the type of distribution.
+Wang et al.
+(2025b) develop a model and framework for modeling and evaluating fine-grained probability estimation of propositions conditioned on context, and show that their model captures some amount of human uncertainty.
+Prior work has also found that several LLMs are capable few-shot regressors, often outperforming traditional supervised methods Vacareanu et al.
+(2024), and capable fine-tuned regressors Song et al.
 (2024).
-Similarly, distribution elicitation methods are techniques used to obtain probability distributions from experts about unknown quantities Falconer et al.
+Distribution prediction in the real world is a complex task, and can rely on both uncertainty quantification and nuanced language understanding Wang et al.
+(2025b); Vedula et al.
 (2025).
+This motivates using LLMs for this task, and prior work has developed methods in various domains for doing LLM-based distribution prediction: combining LLM-generated distributions to better predict human annotator distributions Hashemi et al.
+(2024), developing a quantile regression approach for price prediction Vedula et al.
+(2025), and mapping LLM token predictions to continuous distributions for probabilistic forecasting Gruver et al.
+(2023).
+Distributional alignment
+Recent works have also advocated for a shift to pluralistic alignment due to the ecological fallacy of assuming all members of a group share the same values, beliefs, or opinions Gordon et al.
+(2022); Sorensen et al.
+(2024); Feng et al.
+(2024); Kirk et al.
+(2024); Xie et al.
+(2025).
+Closer to our work, prior works have studied aligning LLM distributions with human response distributions on subjective questions.
+We draw inspiration from Santurkar et al.
+(2023), though we evaluate multiple methods of extracting distributions from LLMs.
+Dominguez-Olmedo et al.
+(2024) study systematic biases with LLM-generated distributions and how this differs from human response distributions for surveys.
+Meister et al.
+(2025) study different distribution elicitation methods for subjective questions.
+Cao et al.
+(2025) fine-tune with a divergence-based loss function.
+Our work introduces the notion of supervised calibration in this space, and studies the effect of alignment methods on individual groups.
 3 Approach
 3.1 Datasets
-We consider three human survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA), and the World Values Survey (WVS).
-Both OQA and WVS have been used to study sociodemographic prompting in prior work Santurkar et al.
-(2023); Durmus et al.
+We consider three survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA) Santurkar et al.
+(2023), and the World Values Survey (WVS).
+OQA and WVS have been used to study alignment in prior work Durmus et al.
+(2024); Masoud et al.
+(2025); AlKhamissi et al.
 (2024).
-These three datasets cover a diverse range of subjective topics, such as perceptions of science and public health, public opinion on gun control, data privacy, and various moral opinions and values.
-To reduce downstream LLM costs, we subset datasets to include all WGM ordinal questions and two questions per category from OQA and WVS.
-This totals 92 questions across WGM (14), OQA (38), and WVS (40).
-Demographics for each dataset are shown in Table 11 (Appendix).
-Over all datasets, questions, and demographics, there are 4,500 human response distributions.
-For each question, model, and distribution elicitation method (Section 3.2), we predict the probability distribution of human responses over answer choices.
-This yields 220,500 generated SD-specific response distributions and 4,500 generated SD-agnostic response distributions.
+These three datasets cover diverse subjective topics, including global perceptions of science and public health (WGM), US public opinion (OQA), and various moral opinions and values (WVS).
+We include all WGM ordinal questions and two questions per category from OQA and WVS, yielding 92 questions across WGM (14), OQA (38), and WVS (40) and a total of 4,500 human response distributions over all datasets, questions, and demographics.
 3.2 Distribution elicitation
 From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group.
-To predict these distributions via LLM, we apply (with some modifications) methods inspired from confidence literature as follows.
-We investigate techniques that work most broadly, supporting use with both black-box and white-box LLMs.
+To generate distributions via LLM, we apply (with some modifications) methods inspired from model confidence literature.
+We investigate techniques that work most broadly, supporting use with open-source, open-model, and API-access LLMs.
 That said, Appendix A.2 reports a smaller scale study with log probabilities.
@@
 Verbalized.
-This is simply directly prompting the model to output a distribution in numbers Geng et al.
-(2024b); Tian et al.
+This is directly prompting the model to output a probability distribution Geng et al.
+(2024); Tian et al.
 (2023).
@@
 (2023).
-In our experiments, we sample times.
+In our experiments, we sample times, following prior work Xiong et al.
+(2024).
 Paraphrase.
@@
 3.3 Prompts
-We use two categories of prompts, and for each category, we prompt with the different elicitation methods. See Appendix B for prompts used.
+We use two categories of prompts, with three variations. For each, we prompt with the 3 different elicitation methods. See Appendix B for the prompts.
 Base. Base prompts exclude any sociodemographic information, asking questions from the original dataset with formatting slightly changed for LLMs. See Appendix Table 14 for full prompts.
@@
 The exact prompt varies with demographic and elicitation method.
-See Appendix Table 13 for full prompts used.
-3.4 Metric
-We use the opinion alignment metric from Santurkar et al.
-(2023) to measure similarity between elicited distributions and ground truth distributions for all results.
-This metric makes use of 1 Wasserstein distance, or earth-mover’s distance.
-As noted by Santurkar et al.
-(2023), this metric takes into account the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon.
-The opinion alignment metric ranges from 0 to 1, but in our results we show the value as a percentage.
+See Appendix Table 15 for full prompts.
+3.4 Metric: Opinion Alignment
+We use the following opinion alignment metric from Santurkar et al. (2023) to measure similarity between an elicited distribution () and ground truth distribution () for the set of questions :
+where refers to Wasserstein distance, or earth-mover’s distance, and is a normalization factor.
+Santurkar et al.
+(2023) note this metric accounts for the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon.
+The opinion alignment metric ranges in , but in our results we show the value as a percentage.
 3.5 Models
@@
 Others are reported in Appendix A.1.
-3.6 Calibration
-We apply supervised regression to transform LLM-generated distributions.
-We split the LLM distributions into train, development, and test sets (60-20-20).
-Results in Section 6 show that far less supervision is typically needed in practice.
-The input to regression is the LLM distribution and the output is the scaled distribution.
-We split the distributions by answer choice before normalizing across choices.
-More formally, for a regression model and LLM distribution parameterized by the values for each answer choice , we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice .
-I.e., we train on pairs of to learn transformed values on held-out test questions.
-We learn a regression model for each dataset-LLM-elicitation setting and do development set hyperparameter tuning and model selection to choose between linear regression (including lasso and ridge) and random forest regression.
 | Model | Base prompt | Sociodemographic prompt | |||||||||||
@@
 | Average | 64.1 | 84.6 | 64.0 | 84.6 | 80.1 | 85.1 | 66.4 | 84.8 | 64.0 | 84.7 | 81.1 | 86.2 | |
+3.6 Calibration
+We apply supervised regression to transform LLM-generated distributions to more closely align them with the ground truth human distributions.
+For example, for the question, How much do you trust doctors and nurses?
+A lot, some, not much, or not at all?, the LLM might generate a distribution such as , while the ground truth human response distribution based on the survey question might be .
+We seek to learn a transformation from the LLM-generated distributions to the human distributions that might be applied to all LLM-generated distributions.
+More formally, for a regression model and LLM distribution parameterized by the values for each answer choice (where is the number of answer choices), we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice .
+We then re-normalize the transformed distribution.
+Training examples.
+The examples for the regression model are the distributions split by answer choices.
+For example, the LLM and human distributions given above, and , would be split into supervised pairs of , , , and .
+We thus train on pairs of to learn transformed values on held-out test questions.
+Regression model.
+We learn a regression model for each dataset-LLM-elicitation-prompt setting (270 regression models for 3 datasets, 15 models, 3 elicitation methods, and 2 prompt categories).
+Hyperparameter tuning and model selection are used to choose between linear regression (including lasso and ridge) and random forest regression for each regressor.
+We use the scikit-learn Pedregosa et al.
+(2011) implementation for each regressor.
+Training.
+The total number of examples for each regression model is x x , where is the class within each sociodemographic group (e.g., female and male for sex), is the question (e.g., how much do you trust doctors?), and is the number of answer choices per question.
+These values are different for each setting (see Table 13 for the sizes of and per dataset).
+We were unable to extract distributions for all questions from some of the smaller LLMs (see Table 17), so the number of training examples differs for them.
+We split the LLM distributions into train, development, and test sets (80%-10%-10%).
+For minimal supervision (Section 6), we train on a random sample of examples ( = [1, 5, 10, 50, 100, 200, 500, 1000, 1200], where each example is the full distribution not split by answer choices) from the training set.
+We repeat this procedure 10 times for different random splits and average.
+Evaluation.
+We optimize the regression models with Mean Squared Error (MSE).
+Since we split distributions during training, we normalize output values so that they sum to 1 before calculating opinion alignment with the ground truth distribution.
+For example, if the output of for the distribution is , we would normalize these values to .
 4 Aligning LLMs with human opinions
-We compare two methods: adding SD information to prompts, and calibrating LLM-generated distributions using some human ground truth data for supervision.
+We compare two methods: adding SD information to prompts and calibrating LLM-generated distributions with human ground truth data for supervision.
 We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods.
 Our main results are shown in Table 1.
-The full results for all models are shown in Appendix Table 5.
-Findings (RQ1, RQ2):
-(1) the effectiveness of baseline SD prompting in generating aligned distributions is model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al.
-(2024); (2) calibration increases opinion alignment in aggregate across models, datasets, and elicitation methods by an average of 16.3%; and (3) calibration reduces variance within and across settings.
+The alignment values for all models are shown in Appendix Table 7 for all the results in this section.
+All significance results are calculated via paired t-test with Bonferroni correction.
 4.1 Does SD prompting improve alignment?
-To study RQ1, we compare SD vs.
-base prompts.
-What is the effect of adding SD information on opinion alignment?
-Note that while prior work studies this question only for majority-voted responses Beck et al.
-(2024); Hu and Collier (2024), we instead look at the effect of adding SD information when evaluating distributions.
+For RQ1, we compare SD vs.
+base prompts to study the effect of adding SD information on opinion alignment.
+Note that while prior work studies this question only for majority-voted responses Hu and Collier (2024); Sun et al.
+(2025a); Mukherjee et al.
+(2024); Masoud et al.
+(2025); AlKhamissi et al.
+(2024), we investigate the effect of adding SD information when comparing distributions.
 We study whether this is consistent across distribution elicitation methods, models, and datasets.
-The results for all models are shown in Appendix Table 5.
-For both base and SD prompts, LLM-generated distributions are generally best with verbalized elicitation (Table 1), though with exceptions (e.g., OLMo-Instruct on the OQA dataset).
-As shown in the table, prompting with SD information does not necessarily increase opinion alignment with human responses.
-In fact, it is often comparable or even lower than prompting without any SD information.
-The degree of effectiveness in creating distributions that align with human opinion appears to be model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al.
-(2024).
+For base and SD prompts, LLM distributions are generally most aligned with verbalized elicitation (Table 1), with exceptions (e.g., on the OQA dataset).
+As shown in the table, prompting with SD information often leads to comparable or even lower opinion alignment than prompting without any SD information, although this is model, dataset, and elicitation method dependent.
 4.2 Can we calibrate LLM distributions?
-We have seen that prompting LLMs with SD information does not consistently increase opinion alignment with human responses.
-Since LLM-generated distributions have fairly low opinion alignment, with RQ2, we question whether we can calibrate these distributions to better align them with human response distributions on these survey datasets.
-The results after regression, compared to pre-regression results, are shown in Table 1, with results for all models in Appendix Table 5.
-Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, and by an average of 16.3%.
-We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment decreases by an average of 4.1%.
-These results suggest that LLM-generated distributions for predicting human responses are somewhat uncalibrated, and a simple supervised regression can lead to higher opinion alignment.
+We have seen that prompting LLMs with SD information does not consistently improve opinion alignment with human responses.
+We next investigate for RQ2 whether calibrating these distributions can more accurately and consistently align them with human response distributions.
+The results after regression, compared to pre-regression results, are shown in Table 1.
+Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, improving by an average of 16.3%, where we average the delta between uncalibrated and calibrated opinion alignment across dataset-LLM-elicitation method settings.
+We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment in these settings decreases by an average of 4.1%.
+These results suggest that LLM-generated distributions are somewhat uncalibrated, and a simple supervised regression can improve alignment.
+| Model | Base prompt | Sociodemographic prompt | |||||||||||||||||
+| WGM | Claude | 66.2 | 65.6 | 85.2 | 59.5 | 60.7 | 85.1 | 89.3 | 88.8 | 87.7 | 65.4 | 61.7 | 84.4 | 61.4 | 58.6 | 84.2 | 89.0 | 79.7 | 89.3 |
+| OLMo | 59.6 | 62.9 | 81.5 | 67.7 | 62.7 | 80.4 | 59.8 | 86.1 | 85.1 | 62.3 | 64.0 | 82.5 | 64.5 | 59.8 | 82.7 | 69.8 | 81.8 | 84.6 | |
+| OQA | Claude | 70.5 | 76.4 | 88.8 | 72.5 | 74.3 | 90.4 | 91.7 | 89.9 | 91.7 | 76.1 | 68.0 | 89.9 | 73.3 | 66.1 | 89.6 | 91.9 | 80.3 | 91.6 |
+| OLMo | 72.6 | 79.3 | 89.0 | 72.3 | 77.0 | 88.4 | 65.4 | 85.2 | 79.9 | 72.8 | 77.1 | 88.4 | 70.5 | 76.3 | 88.6 | 68.5 | 82.9 | 81.5 | |
+| WVS | Claude | 46.5 | 55.5 | 80.3 | 51.0 | 52.8 | 80.3 | 75.2 | 80.3 | 80.3 | 61.0 | 61.4 | 80.4 | 56.8 | 59.8 | 80.4 | 75.6 | 79.7 | 81.7 |
+| OLMo | 75.3 | 76.9 | 80.3 | 74.9 | 73.5 | 80.3 | 86.6 | 86.5 | 86.5 | 58.3 | 72.7 | 79.2 | 60.0 | 69.4 | 77.2 | 86.0 | 85.7 | 89.8 | |
 After calibration, opinion alignment in each setting also has much lower variance.
-Calibrated distributions tend to have a much lower standard deviation; standard deviation is lower in 87.2% of dataset-LLM-elicitation settings, and is on average 1.62 times lower.
-That lower standard deviation leads to higher opinion alignment provides more evidence that LLMs might be exaggerating differences in opinion distributions between demographic groups Cheng et al.
-(2023a, b), and that calibration could be helpful in mitigating this.
-In addition, we find that calibration reduces variance across settings.
+Calibrated distributions have a lower standard deviation in 87.2% of dataset-LLM-elicitation settings, and are 1.62 times lower on average.
+That lower standard deviation leads to higher opinion alignment provides further evidence that LLMs may exaggerate differences in opinion distributions between demographic groups Cheng et al.
+(2023a, b), and that calibration could help mitigate this.
+We also find that calibration reduces variance across settings.
 Figure 2 plots standard deviation vs.
 opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration.
-We see that the standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset.
+Standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset.
+Although there are some settings where uncalibrated distributions have a lower standard deviation than their calibrated counterparts, we emphasize that the variance between settings is much lower after calibration.
 While prior work Beck et al.
-(2024); Hu and Collier (2024) and our own previous results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly.
+(2024); Hu and Collier (2024) and our own results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly.
 This means that with calibration, the choice of any particular model becomes less important.
+We do not find distinct patterns with model size within model family (e.g., Mistral small vs.
+Mistral large).
+Indeed, smaller models are sometimes more aligned than larger models, though across prompting methods, elicitation methods, and datasets, larger models are more aligned more consistently (see Appendix Table 7).
+However, we do find that chain of thought prompting (Section 4.3) appears to increase alignment for smaller models, but not for larger models.
+The smaller Llama models in particular often refused to answer questions (i.e., questions would hit safety restrictions for attempting to role-play as a particular identity group), or would not follow instructions in formatting output distributions.
+We were thus unable to extract distributions for all questions, leading to higher variance in alignment.
+See Appendix Table 17 for percent distributions extracted in each setting.
 | Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
@@
 | All Demographics (WVS) | 75.2 | 80.3 | 64.6 | 80.3 | 72.8 | 80.3 | 86.6 | 86.5 | 73.9 | 77.9 | 74.6 | 81.1 | |
+4.3 The effect of chain of thought prompting
+We compare CoT prompting with standard prompting + calibration for two models, the model that is most aligned on average (Claude 3.5 v2) and a small model (OLMo-2-7B Instruct).
+Our results are shown in Table 2.
+We find that for Claude, CoT does not consistently improve alignment over standard prompting, though it does improve in some settings.
+For base prompting with verbalized elicitation, alignment with CoT is not significantly different than with calibration for 2/3 datasets.
+For SD prompting, calibration creates significantly more aligned distributions for all datasets.
+For OLMo, we find that CoT improves alignment much more consistently, often leading to distributions that are just as aligned as, or better aligned than, calibrated distributions for the verbalized setting.
+Calibration is consistently more aligned than CoT for distributions elicited via paraphrase or self-random.
+Upon qualitative inspection, we found that SD + CoT prompting Claude led to an “overthinking” effect Pu et al.
+(2025), where the model sometimes generated a distribution that exaggerated a particular difference with the ground truth.
+For example, the knowledge that “Men generally report higher confidence in scientific knowledge than women due to societal factors” to generate a distribution for the question “How much do you know about science?” led to the model producing a distribution that was more skewed than the ground truth.
+We hypothesize that this led to lower alignment for Claude (and would be similar for other larger, more capable reasoning models), but not for smaller models like OLMo 7B.
+Future work might study this further.
 5 How does alignment vary across SDs?
-So far, we have studied how well LLMs can predict responses from SD groups aggregated across all groups, with the goal of comparing different models and distribution elicitation methods.
+So far, we have studied alignment with SD groups aggregated across groups, in order to compare different models and distribution elicitation methods.
 However, a particular model or method might be more aligned with some sociodemographics over others.
@@
 We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS).
-We only look at distributions elicited with the verbalized method, as it performed the best with most models across datasets.
+We only look at distributions elicited with the verbalized method, as it leads to the most aligned distributions with most models across datasets.
 We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics.
-We choose 5 out of 15 LLMs; the most recent/powerful from each model family: Claude 3.5 v2, Llama 3.2 90B, Mistral large, OLMo-2 7B Instruct, and Qwen 2.5 72B.
 Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 3.
+We show results for the same 5 LLMs used in Section 4.
 See Appendix A.3 for alignment for other demographics.
-Findings (RQ3):
-(1) Calibration produces distributions more aligned with some demographics over others; (2) Claude models are more highly aligned with OQA dataset SD groups; (3) different models are better aligned with specific demographics.
-5.1 How does calibration affect alignment?
+5.1 The effect of calibration
 First, we compare opinion alignment between base and calibrated distributions.
-We use the regression models trained on all sociodemographic groups and study how such aggregate regression models might calibrate individual sociodemographics.
+We use the regression models trained on all SD groups and study how such aggregate regression models might calibrate individual sociodemographics.
 We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples).
-However, calibrated distributions are more aligned for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions.
-As expected, calibration increases alignment for some demographics and reduces alignment for others.
-We find that it tends to increase alignment of SD groups that were less aligned with base-prompted distributions, such as Central Africa, Central America/Mexico, South America, Southern Africa, and Western Africa for world region.
-Alignment is decreased for very high aligned demographic groups, such as Aus/NZ and Northern Europe with Llama 90B.
-5.2 How does alignment vary across LLMs?
-Next, we compare the opinion alignment between different LLMs.
+However, calibrated distributions are more absolutely aligned than their uncalibrated counterparts for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions.
+This includes all demographic categories, including those in Table 3 and Appendix A.3.
+The delta improvement across demographics is most significant for OLMo on the WGM and OQA datasets.
+As expected, calibration improves alignment for some demographics and reduces alignment for others.
+We find that it tends to improve alignment of SD groups that were less aligned with base-prompted distributions, such as Central America/Mexico, South America, Southern Africa, and Western Africa for world region.
+Alignment decreases for very high aligned demographic groups, such as Aus/NZ and Northern Europe.
+| Model | Base prompt | Sociodemographic prompt | |||||||||||||||||
+| WGM | Claude | 66.2 | 63.5 | 86.4 | 59.5 | 61.7 | 83.7 | 89.3 | 89.0 | 89.2 | 65.4 | 59.8 | 84.1 | 61.4 | 58.1 | 84.2 | 89.0 | 78.9 | 89.0 |
+| Llama | 68.1 | 72.5 | 82.8 | 73.0 | 66.5 | 86.4 | 84.8 | 86.0 | 89.2 | 70.6 | 66.7 | 86.0 | 67.6 | 64.9 | 85.6 | 85.0 | 78.7 | 89.6 | |
+| Mistral | 62.4 | 59.5 | 84.5 | 72.0 | 55.4 | 83.7 | 89.4 | 87.2 | 89.6 | 68.4 | 58.4 | 85.8 | 63.0 | 55.3 | 84.9 | 87.3 | 78.1 | 88.5 | |
+| OLMo | 59.6 | 70.5 | 81.1 | 67.7 | 70.9 | 81.0 | 59.8 | 81.1 | 83.7 | 62.3 | 64.2 | 82.2 | 64.5 | 66.0 | 81.6 | 69.8 | 79.3 | 80.6 | |
+| Qwen | 57.7 | 60.6 | 82.8 | 53.3 | 59.0 | 82.1 | 88.2 | 88.3 | 87.3 | 66.0 | 58.5 | 85.2 | 63.4 | 55.9 | 85.4 | 89.1 | 80.4 | 89.4 | |
+| Average | 62.8 | 65.3 | 83.5 | 65.1 | 62.7 | 83.4 | 82.3 | 86.3 | 87.8 | 66.5 | 61.5 | 84.7 | 64.0 | 60.0 | 84.3 | 84.0 | 79.1 | 87.4 | |
+| OQA | Claude | 70.5 | 77.9 | 88.4 | 72.5 | 75.7 | 89.6 | 91.7 | 91.5 | 91.5 | 76.1 | 68.1 | 89.7 | 73.3 | 66.9 | 89.7 | 91.9 | 79.7 | 92.3 |
+| Llama | 79.3 | 77.6 | 89.7 | 75.3 | 76.6 | 89.6 | 86.8 | 88.8 | 88.6 | 79.2 | 75.0 | 90.0 | 76.1 | 75.5 | 89.8 | 83.4 | 80.5 | 87.1 | |
+| Mistral | 79.5 | 77.4 | 89.6 | 75.3 | 74.0 | 89.0 | 85.0 | 89.7 | 86.9 | 75.8 | 70.5 | 89.8 | 72.4 | 69.9 | 87.5 | 83.8 | 80.7 | 87.8 | |
+| OLMo | 72.6 | 77.0 | 88.9 | 72.3 | 74.6 | 85.9 | 65.4 | 81.2 | 77.4 | 72.8 | 72.3 | 88.6 | 70.5 | 74.4 | 88.4 | 68.5 | 78.8 | 86.6 | |
+| Qwen | 73.9 | 75.9 | 90.1 | 67.0 | 74.4 | 89.3 | 88.4 | 88.9 | 87.2 | 74.4 | 71.4 | 90.2 | 71.3 | 69.3 | 89.3 | 89.2 | 81.0 | 87.8 | |
+| Average | 75.2 | 77.2 | 89.3 | 72.5 | 75.1 | 88.7 | 83.5 | 88.0 | 86.3 | 75.7 | 71.5 | 89.7 | 72.7 | 71.2 | 88.9 | 83.4 | 80.1 | 88.3 | |
+| WVS | Claude | 46.5 | 62.4 | 81.7 | 51.0 | 62.1 | 80.8 | 75.2 | 82.1 | 81.5 | 61.0 | 64.7 | 80.4 | 56.8 | 64.0 | 80.2 | 75.6 | 81.9 | 80.4 |
+| Llama | 61.6 | 70.1 | 79.4 | 59.1 | 68.9 | 80.9 | 64.6 | 79.7 | 83.5 | 62.1 | 70.3 | 80.0 | 59.5 | 72.2 | 80.9 | 67.7 | 78.8 | 86.2 | |
+| Mistral | 48.8 | 63.9 | 81.4 | 44.0 | 61.0 | 80.6 | 72.8 | 81.0 | 81.1 | 54.3 | 65.0 | 82.4 | 51.5 | 62.7 | 81.4 | 76.6 | 78.7 | 82.5 | |
+| OLMo | 75.3 | 72.9 | 80.0 | 74.9 | 74.2 | 80.2 | 86.6 | 84.2 | 87.8 | 58.3 | 68.8 | 78.8 | 60.0 | 72.3 | 77.5 | 86.0 | 78.0 | 89.7 | |
+| Qwen | 39.6 | 58.3 | 82.1 | 42.4 | 56.8 | 81.3 | 74.0 | 78.4 | 81.0 | 49.1 | 59.0 | 81.5 | 49.4 | 56.8 | 81.2 | 73.4 | 78.3 | 76.6 | |
+| Average | 54.4 | 65.5 | 80.9 | 54.3 | 64.6 | 80.8 | 74.6 | 81.1 | 83.0 | 57.0 | 65.6 | 80.6 | 55.4 | 65.6 | 80.2 | 75.9 | 79.1 | 83.1 | |
+5.2 Alignment variation across LLMs
+We now compare opinion alignment between different LLMs.
 For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East.
-Interestingly, Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data.
+Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data.
 For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen.
-In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others.
 On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets.
 This is discussed further in Section 7.
+In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others.
 All models have lowest alignment with middle income populations.
 This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income.
-In other words, income might be less of an important factor in determining the answer to the survey questions for those with middle income.
+In other words, income might be less important in determining survey answers for those with middle income.
 6 Does supervision work with less data?
-Although we used 80% of our data as supervision for our regression models to calibrate the LLM-generated distributions, we study how opinion alignment varies with smaller amounts of supervised data to explore RQ4.
-Findings (RQ4):
-(1) As few as 5 training examples to the regression model can suffice to achieve close to minimal MSE; (2) on average, degradation in alignment for individual demographics is close to zero at 5 examples; (3) demographics with the largest opinion alignment degradation is dependent on base vs.
-SD prompts; and (4) regression models generalize fairly well to unseen datasets.
-6.1 How much supervised data do you need?
-We evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset.
-We average over 10 different random samples for each training data size.
-We find that as few as 5 examples (with 4 answer choices per example) can be enough to achieve close to the minimal MSE for any particular dataset-LLM-elicitation setting.
-We plot Mean Squared Error (MSE) over training data size in Appendix Figure 4 showing that MSE usually converges at 5 examples, though this is model and dataset dependent.
-6.2 How are individual SD groups affected?
-How does using a small set of random examples affect alignment of individual demographics?
+Our results presented thus far used 80% of each dataset for training our supervised regression models to calibrate the LLM-generated distributions.
+We next investigate whether smaller amounts of supervision can suffice for opinion alignment (RQ4).
+6.1 Amount of supervised data required
+As discussed in Section 3.6, we evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset, and average over 10 random samples for each size.
+We find that as few as 1-10 examples (with 4 answer choices per example) can suffice to closely match the minimal MSE for any particular dataset-LLM-elicitation setting.
+Figure 3, plotting Mean Squared Error (MSE) over training data size, shows that MSE usually converges between 1 and 10 examples, though this is model and dataset dependent.
+6.2 Comparison with few-shot prompting
+Given that the regression models appear to converge with so few examples, one might ask whether the regression is actually learning anything useful, or if we could achieve a similar improvement with few shot prompting instead.
+To investigate this, we 5-shot prompt the LLMs with the same examples given to the regression model in the 5 example calibration setting.
+We use a 5-shot prompt since this lies in the middle of how many examples were needed for convergence for the regression models.
+Whereas we averaged over 10 random samples in those experiments, here we use only 3 of the random samples because we found little variation.
+We average alignment over the 3 runs.
+Table 4 compares few shot prompted () distributions to original and calibrated () distributions for base and SD prompts.
+While 5-shot prompted distributions are not as aligned as calibrated distributions in most settings, they are significantly more aligned in a few cases.
+However, calibration leads to distributions that are more aligned more consistently, across models, prompting methods, and distribution elicitation methods.
+6.3 The effect on individual SD groups
+Although we see convergence with fewer examples, does using a small set of random examples affect alignment of individual demographics?
+We calibrate with a set of 5 examples (in the middle of the 1-10 range described above for MSE convergence).
 Although degradation on average is close to zero, individual demographics are affected differently.
-Table 3 shows those most affected and the degradation amount for both base and SD-prompted verbalized distributions.
+Table 5 shows the five SDs most affected and the degradation amount for both base and SD-prompted verbalized distributions.
 We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant.
-Looking at absolute values, using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
+Using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are historically less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
+7 Does post-training impact alignment?
+To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + supervised finetuning (SFT), base + SFT + direct preference optimization (DPO), and the instruct model (base + SFT + DPO + reinforcement learning with verifiable rewards (RLVR)).
+The SFT step involves finetuning on instruction datasets and scaled synthetic data, the DPO step involves preference finetuning with DPO, and the RLVR step involves finetuning targeted at domains where prompts with verifiable rewards can be constructed, such as math OLMo et al.
+(2025).
+We use these four models out-of-the-box via Huggingface and study differences between base vs.
+SD and calibrated vs.
+uncalibrated with the verbalized elicitation method, since it is typically most aligned.
+Figure 4 plots opinion alignment for all OLMo models.
+Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets.
+OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to have less effect in this setting.
+We suspect the increase in alignment on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS.
+The drop in alignment for uncalibrated distributions on WGM and OQA might indicate that OLMo’s post-training data is less aligned with some of the populations of these datasets.
+Overall, calibration leads to more consistent alignment, with and without SD prompting, regardless of post-training or dataset.
+This is in agreement with the rest of our results, where using calibration brings more consistency across settings.
 | Category | Demographic | ||||
@@
 | Employment | Unemployed | 86.9 | 84.9 | 2.0 |
-6.3 Do models generalize out-of-domain?
-We also study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset.
-We train regression models on two of our three datasets and evaluate on the held-out dataset.
-Results are shown in Appendix A.5.
-As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset.
-Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM.
-This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
-7 Does post-training impact alignment?
-To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + SFT, base + SFT + DPO, and the instruct model (base + SFT + DPO + RLVR).
-We study differences between base vs.
-SD and calibrated vs.
-uncalibrated with the verbalized elicitation method, since it generally is best aligned.
-Findings (RQ5):
-Calibrated alignment is similar both before and after each phase of post-training, though pre-calibrated alignment decreases for 2/3 datasets when post-training is added.
-7.1 What is the effect of calibration?
-Figure 3 plots opinion alignment for all OLMo models.
-Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets.
-OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to affect opinion alignment much less.
-7.2 What is the effect of post-training?
-Changes in alignment after various post-training methods is dataset-dependent.
-For WGM, alignment of the base distribution falls significantly after post-training, though alignment for the SD distributions slightly increase with RLVR.
-Calibration alignment stays consistent after post-training.
-On OQA, alignment decreases with SFT, but stays about the same when adding other post-training.
-This is true of base, SD, and calibrated distributions, though calibration alignment decreases the least.
-On WVS however, alignment increases after post-training methods are added, and alignment changes less with calibration.
-We suspect the increase in performance on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS.
-The drop in performance for uncalibrated distributions on WGM and OQA might indicate that OLMO’s post-training data is less aligned with some of the populations of these datasets.
-8 Suggestions for practitioners
-One must acknowledge risks and and exercise caution in using LLMs to approximate responses from diverse SD groups, since fallible LLMs may naturally exhibit random errors and systematic biases in their outputs.
-That said, the ease, speed, and cost of automated approximation may motivate some practitioners in some settings to use LLMs to approximate human responses, provided the predictive accuracy is “good enough” for their needs.
-Whether this is so will depend greatly on the varying accuracy needs of different real-world uses cases, as well as predictive accuracy relative to the SD groups of interest to each use case.
-We also note that such approximation may represent only an initial starting point, providing an initial guess of the data while awaiting for the real human response data to more slowly trickle in from participants.
-In general, we suggest that practitioners: (1) evaluate distributions rather than a single response when evaluating alignment with human responses, and (2) calibrate LLM-generated distributions.
-As our results shown, calibration can be effective with as few as 5 supervised examples, though degradation varies different SD groups, so one must acknowledge the tradeoffs between risk, cost and quality in performing calibration with less supervision.
-We provide a summary of our results for the five models shown in Table 1.
-8.1 Which methods and models work best?
-As shown in Table 1, although the best distribution elicitation method is LLM and dataset dependent, verbalized elicitation leads to the best aligned distributions in the majority of settings and has a higher average across settings.
-Although this is true both pre-calibration and post-calibration, all methods are much more closely aligned post-calibration: without calibration, verbalized elicitation is the most aligned in 90% of settings, and with calibration it is the most aligned in 63% of settings.
-Adding sociodemographic information does not make LLM distributions more aligned with human distributions consistently; SD verbalized distributions are more aligned than their base verbalized counterparts in 67% of settings pre-calibration and 60% of settings post-calibration.
-Overall, using verbalized elicitation (followed by calibration) might be best to obtain the most aligned distributions.
-However, we note that after calibration, the alignment of distributions produced by all the methods are fairly close (within 2% of each other).
-We also find that prompting with SD information is not as important, though might lead to slightly more aligned distributions in aggregate.
-The most aligned models are much more dataset dependent.
-Claude 3.5 v2 is most aligned for WGM and OQA pre-calibration (with and without SD information).
-Post-calibration, Claude is most aligned for OQA, while Llama 3.2 90B is most aligned for WGM.
-For WVS, OLMo 7B Instruct is most aligned (both pre/post calibration).
 8 Discussion
-Essentializing sociodemographics.
-We note that by prompting LLMs with SD information – and evaluating their generated distributions against human participants with those sociodemographics– we are essentializing the sociodemographic to the identities of the survey respondents and their opinions Wang et al.
-(2024).
-However, practitioners still seek to understand opinions of specific populations (e.g., for election polling) where an aligned LLM might be beneficial for faster iteration on survey questions.
-Although sociodemographics are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al.
-(2022); Biester et al.
-(2022); Pei and Jurgens (2023).
+Limitations of demographic modeling.
+Modeling groups wrt.
+their demographics assuredly provides only a narrow, incomplete view of people, and demographics alone do not determine our opinions.
+That said, such models may usefully approximate distributions of group opinions, provided care is exercised.
+We note that our methods for improving alignment should not be viewed as a way to replace human respondents in surveys, but rather a step towards more accurately representing groups.
 The effect of SD prompting.
-While our SD information provides valuable context about us, it alone does not determine our opinions, and it is fallacious to assume otherwise.
-As we have seen, prompting with SD information on its own also does not necessarily lead to more aligned distributions.
+We showed that simply prompting with SD information does not necessarily lead to more aligned distributions.
 Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al.
@@
 One could make personalized predictions for a sample of a group’s individuals members Gordon et al.
+(2022); Hwang et al.
 (2023) rather than a single distributive prediction for the group.
@@
 LLM distributions are uncalibrated.
-We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated – simple regression allows us to scale distributions to be much more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others.
+We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated; simple regression allows us to scale distributions to be more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others.
 We also find that calibration reduces variance, both across examples and across settings.
-Since calibrated distributions generally have higher performance, this might provide evidence for LLMs exaggerating differences between sociodemographic groups; exaggerations that are reduced with calibration.
+Since calibrated distributions generally have higher performance, this suggests LLMs exaggerate differences between sociodemographic groups, which calibration might help mitigate.
 Calibration on individual SD groups.
 We do not find many significant differences between calibrated/uncalibrated distributions at the individual SD group level.
-Calibrating responses from each individual SD would likely improve alignment for that particular SD, which we do not do here.
-However, we caution efficacy may vary by demographic; some demographic traits might be more relevant to the response distributions of certain questions (e.g., urban vs.
+However, we hypothesize that calibrating responses from each individual SD would likely improve its alignment, though we leave this to future work.
+However, efficacy may vary by demographic; some demographic traits are more relevant to the response distributions of certain questions (e.g., urban vs.
 rural areas might affect gun rights questions more than marital status).
-The dataset, LLM, and distribution elicitation method matter.
-As prior work has found when evaluating SD prompting on human majority responses Beck et al.
-(2024); Hu and Collier (2024), in this work we find that opinion alignment is dataset, LLM, and distribution elicitation method dependent.
-However, we find that verbalized elicitation tends to elicit probabilities that are most aligned with human distributions in the most settings, both before and after calibration.
-We find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is dataset-dependent.
+The choice of LLM and distribution elicitation method matters less with calibration.
+We find that distributional opinion alignment is dataset, LLM, and distribution elicitation method dependent, consistent with prior work evaluating majority responses Beck et al.
+(2024).
+However, verbally-elicited distributions are most aligned with human distributions in the most settings.
+We also find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is also dataset-dependent.
+More importantly however, we find that calibration reduces variance in alignment across models and methods; i.e., the choice of model and elicitation method matters less after calibration.
 9 Conclusion
 We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups.
-We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques.
-Our work also offers guidance for those using LLMs to predict human responses in practice, and our benchmark can enable and stimulate future research.
+We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques, as well as increase consistency between settings.
+Our benchmark also serves to help enable and stimulate future research.
 10 Limitations
 Prompting methods.
-We do not study chain of thought or other prompting methods which might increase overall performance.
 We do not experiment with varied temperatures for our generations.
-For elicitation methods, though a larger might improve generated distributions, our choice of is reasonable given inference costs.
+For elicitation methods, though a larger might improve generated distributions, our choice of is reasonable given inference costs and following prior work Xiong et al.
+(2024).
 We use for verbalized elicitation (as opposed to ) since we found little variance between runs.
-We also leave in-context learning or fine-tuning Orlikowski et al.
-(2025) with human responses to future work.
 Logit-based elicitation methods.
-We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs.
-That said, we do also report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.3).
-We find that using log probabilities leads to the most aligned distributions in 10/24 or 42% of settings with calibration.
+We focus on methods that broadly work with open-source, open-model, and API-access LLMs.
+That said, we report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.2).
+We found that log probabilities with calibration yielded the most aligned distributions in 10/24 settings considered.
 Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
-Demographics studied.
-We only studied a subset of the demographics available in the survey datasets, and each demographic individually.
-Future work might study a larger subset of demographics, as well as intersectional demographics.
+Demographics studied. We only studied a subset of the demographics available in the datasets. Future work might study additional and/or intersectional demographics Crenshaw (1991).
 Regression.
 Our regression predicts each answer choice individually and then normalizes the predicted answers.
+We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better.
 We also optimize regression for MSE instead of the alignment metric.
-We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better.
 Future work might also ensemble multiple elicitation methods.
-Ethics Statement
+Ethical Considerations
 In this work, we study how we might align LLMs to the survey response distributions of various social groups.
-While this can be valuable for social good – for practitioners, NLP researchers, or social science researchers seeking to accurately approximate human responses for survey development, annotation tasks, or social science studies – this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
-We also note that evaluating LLM responses against human responses of particular sociodemographic groups often assumes that everyone in a particular demographic group thinks the same way, and essentializes the demographic to an individual’s identity.
-In this work, we try to address this by evaluating against the distribution of human responses for a particular sociodemographic group.
-By doing so, we hope to better align LLMs with various sociodemographic groups in a distributionally pluralistic manner, rather than assume that all members of the group would answer the same way.
+While this can be valuable for creating LLMs that more accurately portray and reason about population groups, this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
+We note that by prompting LLMs with SD information, and evaluating their distributions against humans with those SDs, we are essentializing the SD to the identities of the survey respondents and their opinions Wang et al.
+(2025a).
+Although SDs are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al.
+(2022); Biester et al.
+(2022); Pei and Jurgens (2023).
+By seeking to infer distributions of human responses rather than majority responses for each SD group, we seek to align LLMs with groups in a distributionally pluralistic manner, rather than assume all members of the group would answer similarly.
 Acknowledgments
@@
 This research was supported in part by Cisco, Good Systems222https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), and a grant from Open Philanthropy.
+We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure.
 The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@
 Results are in Table 7.
-A.2 Minimal Supervision
-Results are in Figure 4.
-A.3 Log Probability Results
+A.2 Log Probability Results
 We get the model’s log probabilities for each answer choice Geng et al.
-(2024b) and normalize to get the distribution over all answer choices, following prior work Santurkar et al.
+(2024) and normalize to get the distribution over all answer choices, following prior work Santurkar et al.
 (2023).
@@
 | Std Dev | 13.8 | 0.7 | 12.4 | 1.4 | 7.8 | 2.7 | 9.5 | 0.9 | 9.9 | 1.1 | 6.4 | 2.9 | |
-A.4 Individual Sociodemographic Results
-Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 6, 7, and 8 respectively.
+A.3 Individual Sociodemographic Results
+Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 8, 9, and 10 respectively.
 | Demographic | Claude-3.5-v2 | Llama-3.2-90B | Mistral-large | OLMo-2-7B-I | Qwen-2.5-72B | Average | |||||||
@@
 | Male | 76.3 | 81.2 | 65.9 | 81.2 | 73.9 | 81.2 | 87.8 | 87.7 | 74.8 | 78.8 | 75.7 | 82.0 | |
-A.5 Generalization results
+A.4 Out-of-domain generalization
+We study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset.
+We train regression models on two of our three datasets and evaluate on the held-out dataset.
+Results are shown in Tables 11 and 12.
+As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset.
+Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM.
+This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
 | Model | Base prompt | Sociodemographic prompt | |||||||||||
@@
 Demographics are shown in Table 13.
-| Dataset | Example Questions | Demographics |
-| WGM | How much do you, personally, know about science?
-Do you know a lot, some, not much, or nothing at all?
-| age, education, employment status, gender, income quintile, living area (city or rural), world region |
-| How much do you trust doctors and nurses in your country? | ||
-| OQA | How important, if at all, is being a gun owner to your overall identity? | age, education, income, marital status, political ideology, political party, race, US region, religion, sex |
-| How much confidence, if any, do you have in elected officials to act in the best interests of the public? | ||
-| WVS | Do you agree strongly, agree, disagree, or disagree strongly with the following statement?
-One of my main goals in life has been to make my parents proud.
-| age, education, employment status, household size, immigrant status, income, marital status, religion, sex |
+| Dataset | Qs | Example question | Demographics |
+| WGM | 14 |
+How much do you, personally, know about science? Do you know a lot, some, not much, or nothing at all?
+How much do you trust doctors and nurses in your country? |
+age (15-29, 30-49, 50+), education(primary, secondary, tertiary), employment status (employed full time for an employer, employed full time for self, employed part time but do not want full time, unemployed, employed part time want full time, out of work force), income quintile (poorest 20%, second 20%, middle 20%, fourth 20%, top 20%), living area (city, rural), sex (male, female), world region (Eastern Africa, Central Africa, North Africa, Southern Africa, Western Africa, Central America and Mexico, Northern America, South America, Central Asia, East Asia, Southeast Asia, South Asia, Middle East, Eastern Europe, Northern Europe, Southern Europe, Western Europe, Aus/NZ) |
+| OQA | 38 |
+How important, if at all, is being a gun owner to your overall identity?
+How much confidence, if any, do you have in elected officials to act in the best interests of the public? |
+age (18-29, 30-49, 50-64, 65+), education (Less than high school, High school graduate, Some college, no degree, Associates degree, College graduate/some postgrad, Postgraduate), income (Less than $30,000, $30,000-$50,000, $50,000-$75,000, $75,000-$100,000, $100,000 or more), marital status (Married, Divorced, Separated, Widowed, Never been married), political ideology (Very conservative, Conservative, Moderate, Liberal, Very liberal), political party (Republican, Democrat, Independent, Other), race (White, Black, Asian, Hispanic, Other), US region (Northeast, Midwest, South, West), religion (Protestant, Roman Catholic, Mormon, Orthodox, Jewish, Muslim, Buddhist, Hindu, Atheist, Agnostic, Other, Nothing in particular), sex (Male, Female) |
+| WVS | 40 |
+Do you agree strongly, agree, disagree, or disagree strongly with the following statement? One of my main goals in life has been to make my parents proud.
 Do you agree, disagree or neither agree nor disagree with the following statement? When jobs are scarce, employers should give priority to people of this country over immigrants. |
+age (16-24 years, 25-34 years, 35-44 years, 45-54 years, 55-64 years, 65+ years), education (early childhood, primary, lower secondary, upper secondary, post-secondary, short-cycle tertiary, bachelor or equivalent, master or equivalent, doctoral or equivalent), employment status (full time, part time, self-employed, retired/pensioned, housewife not otherwise employed, student, unemployed), household size (1, 2, 3, 4, 5, 6, 7 persons or more), immigrant status (I am born in this country, I am an immigrant to this country), income (low, middle, high), marital status (Married, Living together as married, Divorced, Separated, Widowed, Single), religion (none, Roman Catholic, Protestant, Orthodox (Russian/Greek/etc.), Jew, Muslim, Hindu, Buddhist, Other Christian), sex (Male, Female) |
 B.2 Base prompts
@@
 Response: |
+B.4 Prompt variations
+Prompt variations are shown in Table 16.
+| Prompt variation | Prompt template |
+| Few shot | Please complete the task below. Follow the examples given. Examples: {5 examples} Task: {Original prompt} |
+| CoT | {Original question} Explain your reasoning step-by-step before answering.
+{Response instructions} Put your final answer at the end of your response between <ANS_START> and </ANS_END>.
+Use a maximum of 500 words for your reasoning and final answer combined.
+|
+Appendix C Extracted distributions
+The percent of distributions extracted from each LLM for each dataset and elicitation method is shown in Table 17.
+| Model | Base prompt | SD prompt | |||||
+| WGM | Claude-3 | 100 | 100 | 100 | 100 | 100 | 100 |
+| Claude-3.5-v1 | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Claude-3.5-v2 | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Llama-3-70B | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Llama-3.1-70B | 100 | 100 | 82.4 | 100 | 100 | 92.8 | |
+| Llama-3.2-1B | 100 | 100 | 0 | 100 | 100 | 1.3 | |
+| Llama-3.2-11B | 100 | 100 | 70.6 | 100 | 100 | 27.1 | |
+| Llama-3.2-90B | 100 | 100 | 100 | 100 | 100 | 94.9 | |
+| Mistral-small | 100 | 100 | 100 | 100 | 100 | 99.7 | |
+| Mistral-large | 100 | 100 | 100 | 100 | 100 | 100 | |
+| OLMo-2-7B-B | 100 | 100 | 47.1 | 100 | 100 | 69.3 | |
+| OLMo-2-7B-S | 100 | 100 | 100 | 100 | 100 | 97.6 | |
+| OLMo-2-7B-D | 100 | 100 | 100 | 100 | 100 | 92.9 | |
+| OLMo-2-7B-I | 100 | 100 | 70.6 | 100 | 100 | 89.1 | |
+| Qwen-2.5-72B | 100 | 100 | 100 | 100 | 100 | 100 | |
+| OQA | Claude-3 | 100 | 100 | 100 | 100 | 100 | 100 |
+| Claude-3.5-v1 | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Claude-3.5-v2 | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Llama-3-70B | 100 | 100 | 91.9 | 100 | 100 | 99.8 | |
+| Llama-3.1-70B | 100 | 100 | 100 | 100 | 12.1 | 93.4 | |
+| Llama-3.2-1B | 100 | 100 | 0 | 100 | 100 | 0.6 | |
+| Llama-3.2-11B | 100 | 100 | 32.4 | 100 | 100 | 7.2 | |
+| Llama-3.2-90B | 100 | 100 | 100 | 100 | 100 | 99.7 | |
+| Mistral-small | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Mistral-large | 100 | 100 | 100 | 100 | 100 | 99.7 | |
+| OLMo-2-7B-B | 100 | 100 | 100 | 100 | 100 | 99.9 | |
+| OLMo-2-7B-S | 100 | 100 | 100 | 100 | 100 | 99.5 | |
+| OLMo-2-7B-D | 100 | 100 | 94.6 | 100 | 100 | 79.8 | |
+| OLMo-2-7B-I | 100 | 100 | 75.7 | 100 | 100 | 79.5 | |
+| Qwen-2.5-72B | 100 | 100 | 100 | 100 | 100 | 99.4 | |
+| WVS | Claude-3 | 100 | 100 | 100 | 100 | 100 | 100 |
+| Claude-3.5-v1 | 100 | 100 | 100 | 19.6 | 19.6 | 19.6 | |
+| Claude-3.5-v2 | 100 | 100 | 100 | 49 | 49 | 49 | |
+| Llama-3-70B | 100 | 100 | 97.9 | 99.4 | 99 | 87.2 | |
+| Llama-3.1-70B | 100 | 100 | 100 | 49 | 49 | 47.2 | |
+| Llama-3.2-1B | 100 | 100 | 0 | 49 | 49 | 0.4 | |
+| Llama-3.2-11B | 100 | 100 | 39.7 | 49 | 49 | 3.7 | |
+| Llama-3.2-90B | 100 | 100 | 100 | 100 | 100 | 98.3 | |
+| Mistral-small | 100 | 100 | 100 | 100 | 100 | 100 | |
+| Mistral-large | 100 | 100 | 100 | 100 | 100 | 99.9 | |
+| OLMo-2-7B-B | 100 | 100 | 100 | 100 | 100 | 99.6 | |
+| OLMo-2-7B-S | 100 | 100 | 100 | 100 | 100 | 99.3 | |
+| OLMo-2-7B-D | 100 | 100 | 76.8 | 100 | 100 | 72.2 | |
+| OLMo-2-7B-I | 100 | 100 | 82.1 | 100 | 100 | 79.3 | |
+| Qwen-2.5-72B | 100 | 100 | 100 | 100 | 100 | 98.6 | |